### v1

# Deep Class-Incremental Learning: A Survey

Abstract
Deep models, e.g., CNNs and Vision Transformers, have achieved impressive achievements in many vision tasks in the closed world. However, novel classes emerge from time to time in our ever-changing world, requiring a learning system to acquire new knowledge continually. For example, a robot needs to understand new instructions, and an opinion monitoring system should analyze emerging topics every day. Class-Incremental Learning (CIL) enables the learner to incorporate the knowledge of new classes incrementally and build a universal classifier among all seen classes. Correspondingly, when directly training the model with new class instances, a fatal problem occurs — the model tends to catastrophically forget the characteristics of former ones, and its performance drastically degrades. There have been numerous efforts to tackle catastrophic forgetting in the machine learning community. In this paper, we survey comprehensively recent advances in deep class-incremental learning and summarize these methods from three aspects, i.e., data-centric, model-centric, and algorithm-centric. We also provide a rigorous and unified evaluation of 16 methods in benchmark image classification tasks to find out the characteristics of different algorithms empirically. Furthermore, we notice that the current comparison protocol ignores the influence of memory budget in model storage, which may result in unfair comparison and biased results. Hence, we advocate fair comparison by aligning the memory budget in evaluation, as well as several memory-agnostic performance measures. The source code to reproduce these evaluations is available at https://github.com/zhoudw-zdw/CIL_Survey/.
Index Terms:
Class-Incremental Learning, Continual Learning, Lifelong Learning, Catastrophic Forgetting
1 Introduction
Recent years have witnessed the rapid progress of deep learning, where deep neural networks have achieved or even surpassed human-level performances in many fields [1, 2, 3]. The typical training process of a deep network requires pre-collected datasets in advance, e.g., large-scale images [4] or texts [5] — the network undergoes training process of the pre-collected dataset multiple epochs. However, training data is often with stream format in the open world [6, 7]. These streaming data cannot be held for long due to storage constraints [8, 9] or privacy issues [10, 11], requiring the model to be updated incrementally with only new class instances. Such requirements trigger the prosperity of the Class-Incremental Learning (CIL) field, aiming to continually build a holistic classifier among all seen classes. The fatal problem in CIL is called catastrophic forgetting, i.e., directly optimizing the network with new classes will erase the knowledge of former ones and result in irreversible performance degradation. Hence, how to effectively resist catastrophic forgetting becomes the core problem in building CIL models.
Figure 1 depicts the typical setting of CIL. Training data emerge sequentially in the stream format. In each timestamp, we can get a new training dataset (denoted as ‘task’ in the figure) and need to update the model with the new classes. For example, the model learns ‘birds’ and ‘dogs’ in the first task, ‘tigers’ and ‘fish’ in the second task, ‘monkeys’ and ‘sheep’ in the third task, etc. Afterward, the model is tested among all seen classes to evaluate whether it has discrimination for them. A good model should strike a balance between depicting the characteristics of new classes and preserving the pattern of formerly learned old classes. This trade-off is also known as the ‘stability-plasticity dilemma’ in neural system [12], where stability denotes the ability to maintain former knowledge and plasticity represents the ability to adapt to new patterns.
Figure 1: The setting of class-incremental learning. Non-overlapping classes arrive sequentially, and the model needs to learn to classify all the classes incrementally. After the learning process of each task, the model is evaluated among all seen classes. An ideal model should perform well in the newly learned classes and remember the former without catastrophic forgetting.
Figure 2: The setting of Class-Incremental Learning (CIL), Task-Incremental Learning (TIL), and Domain-Incremental Learning (DIL). CIL and TIL share the same training protocol, while TIL is much easier during inference, i.e., only requiring classifying among corresponding label spaces. DIL refers to the data stream with distribution change, where new tasks contain the same classes from different domains, e.g., cartoon and clip-art.
Apart from class-incremental learning, there are other fine-grained settings addressing the incremental learning problem, e.g., Task-Incremental Learning (TIL) and Domain-Incremental Learning (DIL). We show these three protocols in Figure 2. TIL is a similar setting to CIL, and both of them observe incoming new classes in new tasks. However, the difference lies in the inference stage, where CIL requires the model to differentiate among all classes. By contrast, TIL only requires classifying the instance among the corresponding task space. In other words, it does not require cross-task discrimination ability. Hence, TIL is easier than CIL, which can be seen as a particular case of CIL. On the other hand, DIL concentrates on the scenario with concept drift or distribution change [13, 14], where new tasks contain instances from different domains but with the same label space. In this case, new domains correspond to the images in clip-art format. In this paper, we concentrate on the CIL setting, which is a more challenging scenario in the open world.
There is also research about CIL before the prosperity of deep learning [15]. Typical methods try to solve the catastrophic forgetting problem with traditional machine learning models. However, most of them address the incremental learning within two tasks, i.e., the model is only updated with a single new stage [16, 17, 18]. Furthermore, the rapid development of data collection and processing requires the model to grasp long-term and massive-scale data streams that traditional machine learning models cannot handle. Correspondingly, deep neural networks with powerful representation ability well suit these requirements. As a result, deep learning-based CIL is becoming a hot topic in the machine learning and computer vision community.
There are several related surveys discussing the incremental learning problem. For example, [11] focuses on the task-incremental learning problem and provides a comprehensive survey. [19] is a related survey on the class-incremental learning field, while it only discusses and evaluates the methods till 2020. However, with the rapid development of the CIL field, many great works are emerging day by day, which substantially boost the performance of benchmark settings [20, 21, 22, 23]. On the other hand, with the prosperity of Vision Transformer (ViT) [24] and pre-trained models, a heated discussion about ViT in CIL is attracting the attention of the community. Other surveys either focuses on the specific field [25, 26, 27] or lacking the performance evolution among state-of-the-arts [28, 29, 30]. Hence, it is urgent to provide an up-to-date survey containing popular methods to speed up the development of the CIL field.
In this paper, we aim for a comprehensive review of deep class-incremental learning methods from three aspects, i.e., data-centric, model-centric, and algorithm-centric. We also provide a holistic comparison among different kinds of methods over benchmark datasets, i.e., CIFAR100 [31] and ImageNet100/1000 [4]. On the other hand, we highlight an important factor in CIL model evaluation, i.e., memory budget, and advocate fair comparison among different methods with an aligned budget. Correspondingly, we holistically evaluate the extensibility of CIL models with budget-agnostic measures. In general, the contribution of this survey can be summarized as follows:
  • We provide a comprehensive survey of deep CIL, including problem definitions, benchmark datasets, and different families of CIL methods. We organize these algorithms taxonomically (Table I) and chronologically (Figure 3) to give a holistic overview of state-of-the-art.
  • We provide a rigorous and unified comparison among different methods on several publicly available datasets, including traditional CNN-backed and modern ViT-backed methods. We also discuss the insights and summarize the common rules to inspire future research.
  • To boost real-world applications, CIL models should be deployed not only on high-performance computers but also on edge devices. Therefore, we advocate evaluating different methods holistically by emphasizing the effect of memory budgets. Correspondingly, we provide a comprehensive evaluation of different methods given specific budgets as well as several new performance measures.
The rest of this paper is organized as follows. First, we give the problem definition of class-incremental learning in Section 2. Afterward, we organize current CIL methods taxonomically in Section 3 and provide comprehensive evaluations in Section 4. Finally, we summarize future directions of class-incremental learning in Section 5 and then conclude this paper.
2 Preliminaries
In this section, we first give the problem formulation of class-incremental learning. We also discuss the other variations and CIL model decomposition. Afterward, we introduce a commonly used auxiliary set, i.e., exemplar set for CIL.
2.1 Problem Formulation
Definition 1.
Class-Incremental Learning aims to learn from an evolutive stream with incoming new classes [32]. Assume there is a sequence of $B$ training tasks [^1] Also known as ‘session,’ ‘phase,’ and ‘stage.’ $\left\{\mathcal{D}^{1},\mathcal{D}^{2},\cdots,\mathcal{D}^{B}\right\}$ without overlapping classes, where $\mathcal{D}^{b}=\left\{\left({\bf x}_{i}^{b},y_{i}^{b}\right)\right\}_{i=1}^{n_{b}}$ is the $b$-th incremental step with $n_{b}$ training instances. ${\bf x}_{i}^{b}\in\mathbb{R}^{D}$ is an instance of class $y_{i}\in Y_{b}$, $Y_{b}$ is the label space of task $b$, where $Y_{b}\cap Y_{b^{\prime}}=\varnothing$ for $b\neq b^{\prime}$. We can only access data from $\mathcal{D}^{b}$ when training task $b$. The ultimate goal of CIL is to continually build a classification model for all classes. In other words, the model should not only acquire the knowledge from the current task $\mathcal{D}^{b}$ but also preserve the knowledge from former tasks. After each task, the trained model is evaluated over all seen classes $\mathcal{Y}_{b}=Y_{1}\cup\cdots Y_{b}$. Formally, CIL aims to fit a model $f({\bf x}):X\rightarrow\mathcal{Y}_{b}$, which minimizes the expected risk:
$f^{*}=\underset{f\in\mathcal{H}}{\operatorname{argmin}}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{t}^{1}\cup\cdots\mathcal{D}_{t}^{b}}\mathbb{I}(y\neq f(\mathbf{x}))\,$, (1)
where $\mathcal{H}$ is the hypothesis space, $\mathbb{I}(\cdot)$ is the indicator function which outputs $1$ if the expression holds and 0 otherwise. $\mathcal{D}_{t}^{b}$ denotes the data distribution of task $b$. A good CIL model satisfying Eq. 1 has discriminability among all classes, which not only works well on new classes but also preserves the knowledge of former ones.
Class Overlapping: Typical CIL setting assumes $Y_{b}\cap Y_{b^{\prime}}=\varnothing$ for $b\neq b^{\prime}$, i.e., there are no overlapping classes in different tasks. However, in the real world, it is common to observe the old classes emerging in new tasks [33]. When $Y_{b}\cap Y_{b^{\prime}}\neq\varnothing$, the setting is called blurry class-incremental learning (Blurry CIL) [34, 35, 36]. It enables the model to revisit former instances in the later stage, which weakens the learning difficulty. Hence, this paper focuses on the typical CIL setting with no overlapping classes.
In the following discussions, we decompose the CIL model into the embedding module and linear layers, i.e., $f({\bf x})=W^{\top}\phi({\bf x})$, [^2] We omit the bias term for ease of discussions. where $\phi(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}^{d}$, $W\in\mathbb{R}^{d\times|\mathcal{Y}_{b}|}$. The linear layer can be further decomposed into the combination of classifiers: $W=[{\bm{w}}_{1},{\bm{w}}_{2},\cdots,{\bm{w}}_{|\mathcal{Y}_{b}|}]$, where ${\bm{w}}_{k}\in\mathbb{R}^{d}$, $|\cdot|$ denotes the size of the set. The logits are then passed to the Softmax activation for further optimization, i.e., the output probability on class $k$ is denoted as:
$\mathcal{S}_{k}(f({\bf x}))=\frac{\exp^{{\bm{w}}_{k}^{\top}\phi({\bf x})/\tau}}{\sum_{j=1}^{|\mathcal{Y}_{b}|}\exp^{{\bm{w}}_{j}^{\top}\phi({\bf x})/\tau}}\,$, (2)
where $\tau$ is the temperature parameter.
Online CIL: Although data comes with stream format in CIL, the model can conduct multi-epoch training with each training task $\mathcal{D}^{b}$, i.e., offline training within each task. There are some works addressing fully online (one-pass) class-incremental learning, where each batch can be processed once and then dropped [37, 38]. It is a specific case of the current setting, and we concentrate on the generalized CIL setting in this paper.
Table I: The taxonomy of class-incremental learning. We classify CIL methods into data-centric, model-centric, and algorithm-centric methods and further decompose them into six sub-categories. The shade color in the last column denotes the subcategory, which is consistent with Figure 3.
Algorithm Category Subcategory Reference
$\S$ 3.1 Data-Centric Class-Incremental Learning $\S\penalty 10000\$3.1.1 Data Replay Direct Replay [39, 35, 40, 41, 42, 43, 44, 45, 46, 47]
Generative Replay [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]
$\S\penalty 10000\$3.1.2: Data Regularization [59, 60, 40, 61, 62, 63]
$\S$ 3.2 Model-Centric Class-Incremental Learning $\S\penalty 10000\$3.2.1 Dynamic Networks Neuron Expansion [64, 65, 66]
Backbone Expansion [67, 68, 69, 70, 71, 20, 21, 22]
Prompt Expansion [23, 72, 73, 74, 75, 76]
$\S\penalty 10000\$3.2.2: Parameter Regularization [77, 78, 39, 79, 80, 81, 82, 83, 84]
$\S$ 3.3 Algorithm-Centric Class-Incremental Learning $\S\penalty 10000\$3.3.1 Knowledge Distillation Logit Distillation [85, 86, 32, 87, 88, 89, 90, 91, 92]
Feature Distillation [93, 94, 95, 96, 97, 98, 99, 100]
Relational Distillation [101, 102, 103, 104, 105]
$\S\penalty 10000\$3.3.2 Model Rectify Feature Rectify [106, 107, 108, 109, 110, 111]
Logit Rectify [93, 112, 87, 113, 114]
Weight Rectify [115, 116, 117]
2.2 Exemplars and Exemplar Set
As defined in Definition 1, in each incremental task, the model is only able to access the current dataset $\mathcal{D}^{b}$. This helps to preserve user privacy and release the storage burden. However, this restriction is relaxed in some cases, and the model can keep a relatively small set, namely exemplar set, to reserve the representative instances from former tasks.
Definition 2.
Exemplar Set [^3] Also known as the ‘replay buffer’ and ‘memory buffer.’ We use them interchangeably in this paper. is an extra collection of instances from former tasks $\mathcal{E}=\left\{\left({\bf x}_{j},y_{j}\right)\right\}_{j=1}^{M}$, $y_{j}\in\mathcal{Y}_{b-1}$. With the help of the exemplar set, the model can utilize $\mathcal{E}\cup\mathcal{D}^{b}$ for the update within each task. The model manages the exemplar set after the training process of each task.
Exemplar Set Management: Since the data stream is evolving, there are two main strategies to manage the exemplar set in CIL [93]. The first way is to keep a fixed number of exemplars per class, e.g., $R$ per class. Under such circumstances, the size of the exemplar set will grow as the data stream evolves — the model keeps $R|\mathcal{Y}_{b}|$ after the $b$-th task. This will result in a linearly growing memory budget, which is inapplicable in real-world learning systems. To this end, another strategy advocates saving a fixed number of exemplars, e.g., $M$. The model keeps $\frac{M}{|\mathcal{Y}_{b}|}]$ instances per class, where $\cdot$] denotes floor function. It helps to keep a fixed size of exemplars in the memory and release the storage burden. In this paper, we use the second strategy to organize the exemplar set.
Exemplar Selection: Exemplars are representative instances of each known class, which need to be selected from the entire training set. An intuitive way to choose the exemplars is to randomly sample exemplars for each class, which results in diverse instances. By contrast, a commonly used strategy is called herding [118, 32], aiming to select the most representative ones of each class. Given the instance set $X=\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{n}\}$ from class $y$, herding first calculates the class center with current embedding $\phi(\cdot)$:
$\mu_{y}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\phi({\bf x}_{i})\,.$ (3)
Afterward, it calculates and ranks the distance of each instance to the class center $\left\|\mu_{y}-\phi({\bf x}_{i})\right\|$ in ascending order. The exemplars are then selected based on the ranking, e.g., the top-$\frac{M}{|\mathcal{Y}_{b}|}]$ instances with the least distance. Since the class center can be seen as the most representative pattern of each class, selecting exemplars near the class center also enhances the representativeness of exemplars. Herding is now a commonly-used strategy to select exemplars in CIL, and we also adopt it in this paper.
3 Class-Incremental Learning: Taxonomy
There are numerous works addressing class-incremental learning in recent years, raising a heated discussion among machine learning and computer vision society. We organize these methods taxonomically from three aspects, i.e., data-centric, model-centric, and algorithm-centric, as shown in Table I. Data-centric methods concentrate on solving CIL with exemplars, which can be further classified into data replay and data regularization. Model-centric methods either regularize the model parameters from drifting away or expand the network structure for stronger representation ability. Lastly, algorithm-centric methods either utilize knowledge distillation to resist forgetting or rectify the bias in the CIL model. We list the representative methods chronologically in Figure 3 to show the research focus of different periods. In the following sections, we will discuss CIL methods from these three aspects.
3.1 Data-Centric Class-Incremental Learning
Facing the notorious catastrophic forgetting, data-centric methods seek help from extra data, e.g., exemplars. An intuitive way is to utilize former data for rehearsal, which enables the model to review former classes and resist forgetting. On the other hand, some works build regularization terms with the extra data, aiming to control the optimization direction to avoid catastrophic forgetting.
3.1.1 Data Replay
‘Replay’ is important in human cognition system [119] — a student facing final exams shall go over the textbooks to recall former memory and knowledge. This phenomenon can also be extended to the network training process, where a network can overcome catastrophic forgetting by revisiting former exemplars. We denote the sequential training process finetuning the model with $\mathcal{D}^{b}$ as ‘Finetune,’ whose loss function can be denoted as:
$\mathcal{L}$ $=\sum_{({\bf x},y)\in\mathcal{D}^{b}}\ell(f({\bf x}),y)\,$, (4)
where $\ell(\cdot,\cdot)$ measures the discrepancy between inputs, e.g., cross-entropy loss. Finetune is also known as the baseline method for class-incremental learning since it only concentrates on learning the new concepts in the current task. Consequently, it suffers severe forgetting since the model pays no attention to going over former ones. To this end, an intuitive way [32] is to save an extra exemplar set $\mathcal{E}$ (as defined in Definition 2) and include it into the model updating process:
$\mathcal{L}$ $=\sum_{({\bf x},y)\in\left(\mathcal{D}^{b}\cup\mathcal{E}\right)}\ell(f({\bf x}),y)\,.$ (5)
Comparing Eq. 5 to Eq. 4, we can find that exemplars are concatenated to the training set for updating, enabling the retrospective review of former knowledge when learning new concepts.
Utilizing the exemplar set for rehearsal is simple yet effective, which leads to numerous following works. The exemplar sampling process is also similar to the active learning protocol [120], and some works propose corresponding sampling measures to select the informative ones. For example, [39] suggests sampling exemplars with high prediction entropy and near the decision boundary. The model will achieve higher generalization ability by replaying these ‘hard’ exemplars. [35] proposes estimating exemplars’ uncertainty via data augmentation. They choose instances with large prediction diversity by aggregating the predictions of multiple augmented instances. Similarly, [40] proposes to sample exemplars with a greedy strategy in online incremental learning. It proves that exemplar selection is equivalent to maximizing the diversity of exemplars with parameters gradient as the feature. If there are no explicit task boundaries, [41, 42] explore the reservoir sampling process to ensure the exemplars are i.i.d. sampled. [43] proposes to formulate the replay process into a bi-level optimization and keep intact predictions on some anchor points of past tasks. [44] introduces data replay in prototypical network [45], and utilizes the exemplars as pseudo-prototypes for embedding evaluation.
Figure 3: The roadmap of class-incremental learning. We organize representative methods chronologically to show the concentration at different stages. Different colors of these methods denote the sub-categories in Table I. Knowledge distillation and data replay dominated the research before the year 2021, while model rectify and dynamic networks became popular after 2021.
Since exemplars are raw images, directly saving a set of instances may consume enormous memory costs. To this end, several works are proposed to build a memory-efficient replay buffer [22]. [46] argues that extracted features $\phi({\bf x})$ are with lower dimension than raw images $\bf x$} and proposes to save features in the exemplar set to release the burden. Similarly, [47] proposes to keep low-fidelity images instead of raw images to save memory budget. However, since the distributions of extracted features and low-fidelity images may differ from the raw images, an extra adaptation process is needed for these methods, adding to the algorithm’s complexity.
The above-mentioned methods achieve competitive performance by replaying former instances in the memory. Moreover, apart from directly saving instances in the exemplar set, generative models show the potential to model the distribution and generate instances [121, 122, 123], which have also been applied to class-incremental learning. We denote the aforementioned methods directly saving instances for replay as ‘direct replay,’ and the methods utilize generative models as ‘generative replay.’
There often exist two models in generative replay-based CIL, i.e., the generative model for data generation and the classification model for prediction. GR [48] firstly proposes to utilize the generative adversarial network (GAN) [121] in CIL. In each updating process, it utilizes the GAN to generate the instances from former classes and then updates GAN and classification model with both old and new classes. ESGR [49] extends GR by saving extra exemplars. It also proposes to train the separate GAN for each incremental task, which does not require updating GAN incrementally. [50] extends GR by introducing the dynamic parameter generator for model adaptation at test time. FearNet [51] explores brain-inspired CIL and uses a dual-memory system in which new memories are consolidated from a network for recent memories. Recently, [52, 53, 54] explore the application of conditional GAN [124] in CIL, and [55, 56] adopt variational auto-encoders (VAE) [122] to model data distribution. Similarly, [57, 58] model each class into a Gaussian distribution and sample instances directly from the class center.
However, the performance of generative replay methods relies on the quality of generated data, which may fail in complex, large-scale inputs [125]. On the other hand, sequentially updating the generative model also suffers the catastrophic forgetting problem [126, 127, 128], which in turn increases the forgetting in incremental model training.
3.1.2 Data Regularization
Apart from directly replaying former data, another group of works tries to regularize the model with former data and control the optimization direction. Since learning new classes will result in the catastrophic forgetting of old ones, the intuitive idea is to make sure that optimizing the model for new classes will not hurt former ones. GEM [59] aims to find the model that satisfies:
$f^{*}=\underset{f\in\mathcal{H}}{\operatorname{argmin}}\sum_{({\bf x},y)\in\mathcal{D}^{b}}\ell\left(f({\bf x}),y\right)\quad$ (6)
$\text{ s.t. }\sum_{({\bf x}_{j},y_{j})\in\mathcal{E}}\ell\left(f({\bf x}_{j}),y_{j}\right)\leq\sum_{({\bf x}_{j},y_{j})\in\mathcal{E}}\ell\left(f^{b-1}({\bf x}_{j}),y_{j}\right)\,$,
where $f^{b-1}$ stands for the incremental model after training the last task $\mathcal{D}^{b-1}$. Eq. 6 optimizes the model with a restriction, which requires the loss calculated with the exemplar set not to exceed the former model. Since exemplars are representative instances from former classes, GEM strikes a balance between learning new classes and preserving former knowledge. Furthermore, it transforms the constraints in Eq. 6 into:
$\left\langle g,g_{old}\right\rangle:=\left\langle\frac{\partial\ell\left(f({\bf x}),y\right)}{\partial\theta},\frac{\partial\ell\left(f,\mathcal{E}\right)}{\partial\theta}\right\rangle\geq 0\,$, (7)
where $g,g_{old}$ denotes the gradients of the current updating step and exemplar set, respectively. Eq. 7 requires the angle between gradients to be acute. If all the inequality constraints are satisfied, then the proposed parameter update is unlikely to increase the loss at previous tasks. However, if violations occur, GEM proposes to project the gradients $g$ to the closest gradient $\tilde{g}$ satisfying the constraints. GEM further transforms the optimization into a Quadratic Program (QP) problem. However, since the regularization in Eq. 7 is defined among all exemplars, it requires calculating loss among exemplar set and solving the QP problem in each optimization step. Hence, optimizing GEM is very time-consuming. To this end, A-GEM [60] is proposed to speed up the optimization by relaxing the constraints in Eq. 7 into a random batch. A similar idea is also adopted in [40].
There are other methods to address the regularization problem with exemplars. For example, Adam-NSCL [61] proposes to sequentially optimize network parameters by projecting the candidate parameter update into the approximate null space of all previous tasks. OWM [62] only allows weights modification in the direction orthogonal to the subspace spanned by all previously learned inputs. LOGD [63] further decomposes the gradients into shared and task-specific ones. In model updating, the gradient should be close to the gradient of the new task, consistent with the gradients shared by all old tasks, and orthogonal to the space spanned by the gradients specific to the old tasks.
3.1.3 Discussions about Data-Centric Methods
Data-centric CIL algorithms mainly concentrate on resisting forgetting with the help of former data. Replay is a simple yet effective strategy, which has been widely applied to image-based camera localization [129], semantic segmentation [130, 131, 132], video classification [133], and action recognition [95]. However, since the exemplar set only saves a tiny portion of the training set, data replay may suffer the overfitting problem [59, 134]. Simultaneously, the data-imbalance problem [135, 136] also occurs due to the gap between the few-shot exemplars and the many-shot training set. Several works address this problem with balanced sampling [21, 112]. On the other hand, data regularization methods rely on specific assumptions, e.g., treating the loss of exemplars as the indicator of forgetting. These assumptions may not stand in some cases [60], which results in poor performance.
Lastly, both of these methods rely on saving exemplars from the history, which may violate user privacy in some cases [137]. Hence, when developing class-incremental learning in privacy-sensitive scenarios, we should consider learning in an exemplar-free [138, 57, 139, 107] manner.
Figure 4: Illustration of network structure evolving in backbone expansion. Left: DER [20] expands a new backbone per incremental task. Middle: FOSTER [21] adds an extra model compression stage, which maintains limited model storage. Right: MEMO [22] decouples the network structure and only expands specialized blocks.
3.2 Model-Centric Class-Incremental Learning
Model-centric CIL methods mainly concentrate on model evolvement in the learning process. For example, dynamic networks assume the network’s capacity is finite and propose expanding the network when needed to enhance representation ability. On the other hand, parameter regularization methods estimate the importance of parameters and regularize important ones to prevent them from drifting away.
3.2.1 Dynamic Networks
Deep neural networks are proven to produce task-specific features [140, 141]. For example, when the training dataset contains ‘cars,’ the model tends to depict the wheels and windows. However, if the model is updated with new classes containing ‘cats,’ the features would be adapted for beards and stripes. Since the capacity of a model is limited, adapting to new features will result in the overwriting of old ones and forgetting [64]. Hence, utilizing the extracted features for beards and strides is inefficient for recognizing a car. To this end, dynamic networks are designed to dynamically adjust the model’s representation ability to fit the evolving data stream. There are several ways to expand representation ability, and we divide them into three sub-groups, i.e., neuron expansion, backbone expansion, and prompt expansion.
Early works focus on neuron expansion, which adds neurons when the representation ability is not enough to capture new classes. DEN [64] formulates the adjusting process into selection, expansion, duplication, and elimination. Facing a new incremental task, the model first selectively retrains the neurons that are relevant to this task. If the retrained loss is still above some threshold, DEN considers expanding new neurons top-down and eliminating the useless ones with group-sparsity regularization. Afterward, it calculates the neuron-wise drift and duplicates neurons that drift too much from the original values. Apart from heuristically expanding and shrinking the network structure, RCL [65] formulates the network expansion process into a reinforcement learning problem and searches for the best neural architecture for each incoming task. Similarly, Neural Architecture Search (NAS) [142] is also adopted to find the optimal structure for each of the sequential tasks [66].
Expanding neurons shows competitive results with expandable representation. Correspondingly, several works try to duplicate the backbone network for stronger representation ability. PNN [67] proposes learning a new backbone for each new task and fixing the former in incremental learning. It also adds layer-wise connections between old and new models to reuse former features. Expert Gate [68] also expands the backbone per incremental task, while it requires learning an extra gate to map the instance to the most suitable pathway during inference. To release the expansion cost, P&C [69] suggests a progression-compression protocol — it first expands the network to learn representative representations. Afterward, a compression process is conducted to control the total budget. [70, 71] maintain a dual-branch network for class-incremental learning, one for fast adaptation and one for slow adaption. Recently, DER [20] has been proposed to address the CIL problem. Similar to PNN, it expands a new backbone when facing new tasks and aggregates the features with a larger FC layer. Take the second incremental task for an example, where the model output is aggregated as:
$f({\bf x})=W_{new}^{\top}[\phi_{old}({\bf x}),\phi_{new}({\bf x})]\,$, (8)
where $\phi_{old}$ is the former backbone, and $\phi_{new}$ is the newly initialized backbone. $\cdot$$\cdot$] denotes feature aggregation, and $W_{new}\in\mathbb{R}^{2d\times|\mathcal{Y}_{b}|}$ is the newly initialized FC layer. In the model updating process, the old backbone is frozen to maintain former knowledge:
$\mathcal{L}=\sum_{k=1}^{|\mathcal{Y}_{b}|}-\mathbb{I}(y=k)\log\mathcal{S}_{k}(W_{new}^{\top}[\bar{\phi}_{old}({\bf x}),\phi_{new}({\bf x})])\,$, (9)
where $\bar{\phi}_{old}({\bf x})$ denotes the old backbone is frozen. DER also adopts an auxiliary loss to differentiate between old and new classes. Eq. 9 depicts a way to continually expand the model with new features. Under such circumstances, if the old backbone is optimized with ‘cars,’ the features extracted by $\phi_{old}$ are then representative of wheels and windows. The new backbone trained with ‘cats’ is responsible for extracting beards and stripes. Since the old backbone is frozen in later stages, learning new classes will not overwrite the features of old ones, and forgetting is alleviated. Figure 4 (left) depicts the model evolution of DER.
DER has achieved powerful performance with humble forgetting. However, saving a backbone per task requires numerous memory size, and many works are proposed to obtain expandable features with a limited memory budget. FOSTER [21] formulates the learning process in Eq. 9 as a feature-boosting [143] problem. It argues that not all expanded features are needed for incremental learning and need to be integrated to reduce redundancy. For example, suppose old classes contain ‘tigers,’ and new classes contain ‘zebras.’ In that case, the stripe will be a useful feature that could be extracted by both old and new backbones. Under such circumstances, forcing the new backbone to extract the same kind of features is less effective for recognition. Hence, FOSTER adds an extra model compression process by knowledge distillation [144]:
$\min_{f_{s}({\bf x})}\operatorname{KL}\left(\mathcal{S}\left(f_{t}({\bf x})\right)\|\mathcal{S}\left(f_{s}({\bf x})\right)\right)\,.$ (10)
Eq. 10 aims to find the student model $f_{s}$ that has the same discrimination ability as the teacher model $f_{t}$ by minimizing the discrepancy between them. The teacher is the frozen expanded model with two backbones: $f_{t}({\bf x})=W_{new}^{\top}[\phi_{old}({\bf x}),\phi_{new}({\bf x})]$, and the student is the newly initialized model $f_{s}({\bf x})=W^{\top}\phi({\bf x})$. Hence, the number of backbones is consistently limited to a single one, and the memory budget will not suffer catastrophic expansion. Figure 4 (middle) depicts the model evolution of FOSTER.
Recently, MEMO [22] has been proposed to address the memory problem in CIL, aiming to enable model expansion with the least budget cost. It finds that in class-incremental learning, shallow layers of different models are similar, while deep layers are diverse. In other words, shallow layers are more generalizable, while deep layers are specific to the task, making expanding shallow layers less memory-efficient for CIL. Hence, MEMO proposes to decouple the backbone at middle layers: $\phi({\bf x})=\phi_{s}(\phi_{g}({\bf x}))$, where specialized block $\phi_{s}$ corresponds to the deep layers in the network, while generalized block $\phi_{g}$ corresponds to the rest shallow layers. Compared to DER, MEMO only expands specialized blocks $\phi_{s}$, and transforms Eq. 9 into:
$\sum_{k=1}^{|\mathcal{Y}_{b}|}-\mathbb{I}(y=k)\log\mathcal{S}_{k}(W_{new}^{\top}[{\phi_{s}}_{old}(\phi_{g}({\bf x})),{\phi_{s}}_{new}(\phi_{g}({\bf x}))])\,$,
which indicates that task-specific deep layers can be built for each task upon the shared shallow layers $\phi_{g}({\bf x})$. Figure 4 (right) depicts the model evolution of MEMO.
Recently, Vision Transformer (ViT) [24] has attracted the attention of the computer vision community, and many works tend to design CIL learners using ViT as the backbone. DyTox [23] is the first work to explore ViT in CIL, which finds that model expansion in ViT is much easier than in convolutional networks. In DyTox, only task tokens are expanded for each new task, which requires much less memory than saving the whole backbone. Similarly, L2P [72] and DualPrompt [73] explore how to build CIL learners with pre-trained ViT. They borrow ideas from Visual Prompt Learning (VPT) [145] to incrementally finetune the model with prompts. In L2P, the pre-trained ViT is frozen during the learning process, and the model only optimizes the prompts to fit new patterns. The prompt pool is defined as:
$\mathbf{P}=\left\{P_{1},P_{2},\cdots,P_{M}\right\}\,$, (11)
where $M$ is the total number of prompts, $P_{i}\in\mathbb{R}^{L_{p}\times d}$ is a single prompt with token length $L_{p}$ and the same embedding size $d$ as the instance embedding $\phi({\bf x})$. The prompts are organized as key-value pairs — each instance selects the most similar prompts in the prompt pool via KNN search. It obtains instance-specific predictions by adapting the input embeddings as:
${{\bf x}}_{p}=\left[P_{s_{1}};\cdots;P_{s_{N}};\phi({\bf x})\right],\quad 1\leq N\leq M\,.$ (12)
$P_{s_{j}}$ are selected prompts via KNN search. The adapted embeddings are then fed into the self-attention layers [146] to obtain instance-specific representations. CODA-Prompt [74] extends the prompt search with the attention mechanism. Apart from pre-trained ViT, S-Prompt [75] utilizes the pre-trained language-vision model CLIP [147] for incremental learning, which simultaneously learns language prompts and visual prompts to boost representative embeddings. [76] extends CLIP with spatial and temporal prompts for incremental video classification.
3.2.2 Parameter Regularization
Dynamic networks seek to adjust model capacity with data evolves. However, if the model structure is fixed and unchangeable, how to adjust the plasticity to resist catastrophic forgetting? Parameter regularization methods consider that the contribution of each parameter to the task is not equal. Hence, they seek to evaluate the importance of each parameter to the network and keep the important ones static to maintain former knowledge.
Typical works estimate a distribution over the model parameters and use it as the prior when learning new tasks. Due to the large amounts of parameters, the estimation process often assumes them to be independent. EWC [77] is the first work addressing parameter regularization. It maintains an importance matrix with the same scale of the network, i.e., $\Omega$. Denote the $k$-th model parameters as $\theta_{k}$, the importance of $\theta_{k}$ is represented by $\Omega_{k}\geq 0$ (the larger $\Omega_{k}$ indicates $\theta_{k}$ is more important). Apart from the training loss in Eq. 4 to learn new classes, EWC builds an additional regularization term to remember old ones:
$\mathcal{L}=\ell(f({\bf x}),y)+\frac{1}{2}\lambda\sum_{k}\Omega_{k}(\theta_{k}^{b-1}-\theta_{k})^{2}\,.$ (13)
The parameter-wise regularization term is calculated based on two parts. $\theta_{k}^{b-1}$ denotes the $k$-th parameter after learning last task $\mathcal{D}^{b-1}$. Hence, $(\theta_{k}^{b-1}-\theta_{k})^{2}$ represents the parameter drift from the last stage, and $\Omega_{k}$ weighs it to make sure important parameters do not shift away from the last stage. Since the model at the last stage represents the ‘old’ knowledge, consolidating important parameters can fix the knowledge from being forgotten.
Eq. 13 depicts a way to penalize essential parameters, and there are different ways to calculate the importance matrix $\Omega$. In EWC, Fisher information matrix [148] is adopted to estimate $\Omega$. However, the importance calculation in EWC is conducted at the end of each task, which ignores the optimization dynamics along the model training trajectory. To this end, SI [78] proposes to estimate $\Omega$ in an online manner and weigh the importance via its contribution to loss decay. RWalk [39] combines these importance estimation techniques. [79, 80] resort to an extra unlabeled dataset for online evaluation. IMM [81] finds a maximum of the mixture of Gaussian posteriors with the estimated Fisher information matrix. IADM [82] and CE-IDM [83] analyze the capacity and sustainability of different layers in class-incremental learning and find that different layers have different characteristics in CIL. In detail, shallow layers converge faster but have limited representation ability. By contrast, deep layers converge slowly while having powerful discrimination abilities. Hence, IADM augments EWC with an ensemble of different layers and learns layer-wise importance matrix in an online manner. K-FAC [84] extends the Fisher information matrix approximation with the Kronecker factorization technique.
3.2.3 Discussions about Model-Centric Methods
Model-centric methods either design model structure adjustment techniques or build generalization terms to resist forgetting. Apart from these groups, there are works addressing network masks to divide a large network into sub-networks for each task [149, 150, 151, 152]. However, deciding the activation of a specific sub-network requires the task identifier or learning extra task classifiers. On the other hand, several works [153, 154, 155] propose to design specific modules for incremental new tasks, e.g., adapters [156]. However, manually handcrafting these modules requires heuristic designs or task-specific priors.
Learning dynamic networks, especially backbone expansion methods, have achieved state-of-the-art performance in recent years [22]. However, it often requires expandable memory budgets, which is unsuitable for incremental learning on edge devices. Secondly, pre-trained models are needed for the prompt expansion methods. However, a pre-trained model is not always available for some specific downstream tasks, e.g., face recognition [157], speech recognition [158], etc. Therefore, how to get rid of the dependence on pre-trained models is essential for these methods in real-world applications.
On the other hand, parameter regularization methods have attracted the attention of the community in the early years. However, estimating parameter importance requires saving the matrix with the same scale as the model. It faces the same risk as dynamic networks that the memory budget is linearly increasing. On the other hand, the importance matrix may conflict at different incremental stages [92], making it hard to optimize the model.
Figure 5: Illustration of knowledge distillation methods in class-incremental learning. Left: Logit distillation (e.g., iCaRL [32]) aligns the model outputs to make the old model and new model share the same semantic relationship. Middle: Feature distillation (e.g., UCIR [93]) aligns the features produced by the old and new models to make sure the new model does not forget old features. Right: Relational distillation (e.g., R-DFCIL [101]) resorts to structural inputs, e.g., triples, and aligns the input relationship of the old and new model.
3.3 Algorithm-Centric Class-Incremental Learning
Algorithm-centric CIL methods focus on designing algorithms to maintain the model’s knowledge in former tasks. For example, the core idea of knowledge distillation-based CIL methods is to build the mapping between old and new models and reflect the characteristics of the old model in the updating process. On the other hand, model rectify-based methods aim to discover and reduce the bias in incremental models.
3.3.1 Knowledge Distillation
Training data is evolving in the incremental learning process, requiring tuning the model sequentially. We can denote the model after the previous stage $f^{b-1}$ as the ‘old model’ and the current updating model $f$ as the ‘new model.’ Assuming the old model is a good classifier for all the seen classes in $\mathcal{Y}_{b-1}$, how can we utilize it to resist forgetting in the new model?
To enable the old model to assist the new model, an intuitive way utilizes the concept proposed in [159, 160, 144], i.e., knowledge distillation (KD). KD enables the knowledge transfer from a teacher model to the student model, with which we can teach the new model not to forget. There are several ways to build the distillation relationship, and we divide these KD-based methods into three subgroups, i.e., logit distillation, feature distillation, and relational distillation.
LwF [85, 86] is the first success to apply knowledge distillation into CIL. Similar to Eq. 13, it builds the regularization term via knowledge distillation to resist forgetting:
[…]
where the old model ${f}^{b-1}$ is frozen during updating. The regularization term builds the mapping between the old and new models by forcing the predicted probability among old classes to be the same. Given a specific input $\bf x$}, the output probability of the $k$-th class reveals the semantic similarity of the input to this class. Hence, Eq. 14 forces the semantic relationship of the old model and new model to be the same and resists forgetting. iCaRL [32] extends LwF with the exemplar set, which helps to further recall former knowledge during incremental learning. Additionally, it drops the fully-connected layers and follows [161] to utilize nearest-mean-of-exemplars during inference. Eq. 14 strikes a trade-off between old and new classes, where the former part aims to learn new classes and the latter one maintains old knowledge. Since the number of old and new classes may differ in different incremental stages, BiC [87] extends Eq 14 by introducing a dynamic trade-off term:
$\mathcal{L}=(1-\lambda)\ell(f({\bf x}),y)+\lambda\sum_{k=1}^{|\mathcal{Y}_{b-1}|}-\mathcal{S}_{k}({f}^{b-1}({\bf x}))\log\mathcal{S}_{k}(f({\bf x}))\,$,
where $\lambda=\frac{\lvert\mathcal{Y}_{b-1}\rvert}{\lvert\mathcal{Y}_{b}\rvert}$ denotes the proportion of old classes among all classes. It increases as incremental tasks evolve, indicating that the model pays more attention to old ones.
LwF inspires the community to build the mapping between models, making knowledge distillation a useful tool in CIL. D+R [88] suggests changing the first part in Eq. 14 into a distillation loss by training an extra expert model. GD [89] proposes to select wild data for model distillation and designs a confidence-based sampling method to effectively leverage external data. Similarly, DMC [90] proposes to train a new model in each incremental stage and then compress them into a single model via an extra unlabeled dataset. Finally, if no additional data is available for knowledge distillation, ABD [91] proposes distilling synthetic data for incremental learning. These methods only concentrate on utilizing the old model to help resist forgetting in the new model. However, COIL [92] suggests conducting bidirectional distillation with co-transport, where semantic relationships between old and new models are both utilized.
Apart from distilling the logits, some works propose to distill the intermediate product of deep models, e.g., extracted features. UCIR [93] replaces the regularization term in Eq. 14 into:
$\mathcal{L}=\ell(f({\bf x}),y)+(1-\langle\frac{\phi^{b-1}({\bf x})}{\lVert\phi^{b-1}({\bf x})\rVert},\frac{\phi({\bf x})}{\lVert\phi({\bf x})\rVert}\rangle)\,.$ (15)
Eq. 15 forces the features extracted by the new embedding module to be the same as the old one, which is a stronger regularization than Eq. 14. Several works follow it to utilize feature distillation in CIL [94, 95], while others address distilling other products. LwM [96] suggests penalizing the changes in classifiers’ attention maps to resist forgetting. AFC [97] conducts the distillation considering the importance of different feature maps. PODNet [98] minimizes the difference of the pooled intermediate features in the height and width directions instead of performing element-wise distillations. CVIC [162] decouples the distillation term into spatial and temporal features for video classification. DDE [99] distills the causal effect from the old training to preserve the old knowledge. GeoDL [100] proposes to conduct distillation based on the projection of two sets of features from old and new models.
However, both logit and feature distillation address the instance-wise mapping between old and new models. To reveal the structural information in model distillation, several works suggest conducting relational knowledge distillation [163]. The differences between these groups of knowledge distillation are shown in Figure 5.
To conduct relational distillation, a group of instances needs to be extracted, e.g., triplets. We denote the extracted triplets as $\{{\bf x}_{i},{\bf x}_{j},{\bf x}_{k}\}$, where ${\bf x}_{i}$ is called the anchor. In a triplet, the target neighbor ${\bf x}_{j}$ is similar to the anchor ${\bf x}_{i}$ with the same class, while the impostor ${\bf x}_{k}$ is dissimilar to ${\bf x}_{i}$ (usually from different classes). R-DFCIL [101] suggests mapping the angle among triplets:
$\sum_{\{{\bf x}_{i},{\bf x}_{j},{\bf x}_{k}\}\in\mathcal{D}^{b}}\left\|\cos\angle\mathbf{t}_{i}\mathbf{t}_{j}\mathbf{t}_{k}-\cos\angle\mathbf{s}_{i}\mathbf{s}_{j}\mathbf{s}_{k}\right\|\,$, (16)
where $\mathbf{t}_{m}=\phi^{b-1}({\bf x}_{m})$ is the representation in the old model’s embedding space, and $\mathbf{s}_{m}=\phi({\bf x}_{m})$ denotes the representation in the current model. The cosine value is calculated in the corresponding embedding space. Eq. 16 provides a way to encode the old model’s structural information into the new model and align the feature space softly. ERL [102] extends this regularization into few-shot CIL scenarios. TPCIL [103] models the relationship with the elastic Hebbian graph and penalizes the changing of the topological relations between vertices. TOPIC [104] further explores neural gas network to model the class-wise relationship. Apart from the triplet relationship, MBP [105] extends the regularization to the instance neighborhood, requiring the old and new models to have the same distance ranking in the neighborhood.
3.3.2 Model Rectify
Assuming we can get all the training datasets at once and shuffle them for training with multiple epochs, the model will not suffer any forgetting and will perform well among all classes. Such a protocol is known as the upper bound of class-incremental learning, denoted as ‘Oracle.’ However, since the models trained with incremental data suffer catastrophic forgetting, several methods try to find the abnormal behaviors in CIL models and rectify them like the oracle model. These abnormal behaviors include the output logits, classifier weights, and feature embedding.
The first method addressing the bias of the CIL model is UCIR [93]. It finds that the weight norm of new classes is significantly larger than old ones, and the model tends to predict instances as the new classes with larger weights. Hence, UCIR proposes to utilize a cosine classifier to avoid the influence of biased classifiers: $f({\bf x})=\frac{W}{\lVert W\rVert}^{\top}\frac{\phi({\bf x})}{\lVert\phi({\bf x})\rVert}$. Hence, the weight norm will not influence model predictions in incremental learning. WA [115] further normalizes the weight after every optimization step. It also introduces weight clipping to ensure the predicted probability is proportionate to classifier weights. SS-IL [116] explains the reason for weight drifting, which is caused by the imbalance phenomena between old and new instances. Since the number of new class instances is much more than that of old ones, optimizing the model with cross-entropy loss will increase the weight of new classes and decrease old ones. Hence, SS-IL suggests separated softmax operation and task-wise knowledge distillation to alleviate the influence of imbalanced data. RPC [117] claims that the classifiers of all classes can be pre-allocated and fixed. This makes it impossible for classifiers to be biased toward new classes.
On the other hand, several works find that the predicted logits of new classes are much larger than old ones. E2E [112] proposes to finetune the fully-connected layers with a balanced dataset after each stage. Furthermore, BiC [87] proposes to attach an extra rectification layer to adjust the predictions. The extra layer only have two parameters, i.e., re-scale parameter $\alpha$ and bias parameter $\beta$, and the rectified output for the $k$-th class is denoted as:
$\hat{f}({\bf x})_{k}=\begin{cases}\alpha{\bm{w}}_{k}^{\top}\phi({\bf x})+\beta,&k\in{Y}_{b}\\ {\bm{w}}_{k}^{\top}\phi({\bf x}),&otherwise\end{cases}\,.$ (17)
Note that only the logits for new classes ($k\in{Y}_{b}$) are rectified after each incremental task. BiC separates an extra validation set from the exemplar set, i.e., $\mathcal{E}=\mathcal{E}_{train}\cup\mathcal{E}_{val}$, and uses the validation set to tune the rectification layer. On the other hand, IL2M [113] suggests re-scaling the outputs with historical statistics. Suppose an instance is predicted as a new class. In that case, the logits will be re-scaled to ensure the predictions of old and new classes follow the same distribution. AANets [114] learns to adaptively aggregate the stable and plastic blocks and adjust the predictions.
Lastly, since the embedding module is sequentially updated in CIL, some works try to rectify the biased representations of incremental models. For example, SDC [107] utilizes a nearest-mean-of-exemplars classifier, which calculates the class centers and assigns instances to the nearest class center. However, since the embedding is incrementally updated, the class centers calculated in the former stage may suffer a drift in the next stage, making the classification results unreliable. Since old class instances are not available in the current stage, SDC aims to calibrate the class centers of old classes with the drift of new classes. CwD [108] analyzes the differences of embeddings among the CIL model and oracle model and finds that the embeddings of the oracle model scatter more uniformly. It aims to make the CIL model similar to the oracle by enforcing the eigenvalues to be close. ConFiT [109] relieves the feature drift from the middle layers. Other works address the model weight rectification. CCLL [110] aims to calibrate the activation maps of old models during incremental learning. RKR [111] proposes to rectify the convolutional weights of old models when learning new tasks. Recently, FACT [106] depicts a new training paradigm for CIL, namely forward compatible training. Since the embedding space is endlessly adjusted for new classes, FACT proposes to pre-assign the embedding space for new classes to relieve the burden of embedding tuning.
3.3.3 Discussions about Algorithm-Centric Methods
Algorithm-centric methods design task-specific algorithms to resist forgetting. Knowledge distillation-based methods have been widely applied to various incremental tasks, e.g., semantic segmentation [164, 165, 166, 167], person re-identification [168, 169], human action recognition [95], and federated learning [137]. However, since the knowledge distillation term aims to strike a balance between learning the new and remembering the old, it will undoubtedly suffer catastrophic forgetting compared to dynamic networks. [22] finds that knowledge distillation-based methods and dynamic networks have their advantages given different memory budgets. Specifically, knowledge distillation methods show stronger performance given limited memory, while dynamic networks require an adequate memory budget to perform competitively.
On the other hand, model rectification methods aim to reduce the biased inductive bias in the CIL model and align it to the oracle model. This line of work helps to understand the inherent factor of catastrophic forgetting. Apart from the rectification methods listed in Section 3.3.2, some methods [170, 116] address that the bias of the CIL model is from the imbalanced data stream. [171] finds that the embedding trained with contrastive loss suffers less forgetting than cross-entropy loss. [172] finds that the batch normalization layers [173] are biased in CIL and proposes to re-normalize the layer outputs. It is worth exploring other factors that could result in catastrophic forgetting and corresponding solutions in the future.
3.4 Discussions about Classification Criterion and Scopes of This Survey
In this paper, we divide current class-incremental learning methods into three groups based on the key point of these methods. For the methods discussing how to save and utilize exemplars, we divide them into data-centric methods. Similarly, for the methods designing model expansion and parameter-wise regularization, we denote them as model-centric methods. Lastly, we assign the methods that develop specific regularization to pass knowledge from the old to the new model and rectify the current model to the algorithm-centric group. It must be noted that these methods also learn from each other, and there is no strict boundary between them. For example, iCaRL [32] utilizes the knowledge distillation and exemplar replay, but we assign it to the knowledge distillation-based methods. In fact, data replay is becoming a standard protocol in the class-incremental learning community, which is widely adopted in most methods. As shown in Figure 3, dynamic networks and knowledge distillation methods have dominated publications in recent years. However, it does not indicate that traditional methods like data replay are no longer popular. By contrast, most of these methods rely on the exemplar set to get better performance.
On the other hand, in the discussion of this paper, we mainly focus on the development of class-incremental learning in the image classification field. The main reason is that most influential works about CIL address the image classification problem. Furthermore, it has shown the potential to apply the techniques in classification into other tasks, e.g., semantic segmentation [164, 165, 166, 167, 174], object detection [175, 176, 177, 178, 179, 180, 181], video understanding [95, 133], medical surgery [182, 183], language model [184, 185, 186, 187], language-vision model [188, 189], etc. Hence, the scope of our discussion is mainly focused on image classification tasks.
4 Experimental Evaluation
In this section, we conduct comprehensive experiments to evaluate the performance of different kinds of CIL methods with benchmark datasets. We first introduce the benchmark experimental setting, dataset split, evaluation protocol, and implementation details. Afterward, we aim to compare these methods from three aspects:
  • How do these methods perform on benchmark datasets?
  • Are they fairly compared? How to fairly compare them?
  • How to evaluate them with memory-agnostic measure?
Sections 4.2 and 4.3 answer the first question, and Section 4.5 answers the second question. We provide the holistic performance measures in Section 4.6 to answer the third question. We also analyze the influence of incremental learning by visualizing the confusion matrix in Section 4.4 and summarize the comparison results in Section 4.7. Finally, we report more experimental results, implementation details, and visualizations of fully connected layers in the supplementary material.
(a) CIFAR100 Base0 Inc5
(b) CIFAR100 Base0 Inc10
(c) CIFAR100 Base50 Inc10
(d) CIFAR100 Base50 Inc25
Figure 6: Incremental accuracy of different methods on CIFAR100. Legends are shown in (d).
4.1 Experimental Settings
4.1.1 Benchmark Datasets
iCaRL [32] firstly formulates the comparison protocol of class-incremental learning, which was widely followed and compared in other works. It suggests using CIFAR100 [31] and ImageNet100/1000 [190] for evaluation. They are listed as:
  • CIFAR100: There are 100 classes with 60,000 images, in which 50,000 are training instances and 10,000 are testing ones, with 100 images per class. Each image is represented by $\times32$ pixels.
  • ImageNet1000: is a large-scale dataset with 1,000 classes, with about 1.28 million images for training and 50,000 for validation. The spatial size of images varies.
  • ImageNet100: is the subset of ImageNet1000 containing 100 classes [32]. These classes are selected from the first 100 classes after a random shuffle.
There are other works [77, 75, 116, 92, 104, 139, 82] utilizing MNIST [191], Core50 [192], Google Landmark-v2 [193], CUB200-2011 [194], TinyImageNet [195], and miniImageNet [190] for model evaluation, while the aforementioned datasets are the most widely adopted in the current CIL community. Hence, we choose CIFAR100 and ImageNet100/1000 for model evaluation.
Table II: Average and last accuracy performance comparison on CIFAR100. ‘#P’ represents the number of parameters (million). The full table is reported in the supplementary.
Method Base0 Inc5 Base0 Inc10
#P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$} #P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$}
Finetune 0.46 17.59 4.83 0.46 26.25 9.09
EWC 0.46 18.42 5.58 0.46 29.73 12.44
LwF 0.46 30.93 12.60 0.46 43.56 23.25
GEM 0.46 31.73 19.48 0.46 40.18 23.03
Replay 0.46 58.20 38.69 0.46 59.31 41.01
RMM 0.46 65.72 51.10 0.46 68.54 56.64
iCaRL 0.46 63.51 45.12 0.46 64.42 49.52
PODNet 0.46 47.88 27.99 0.46 55.22 36.78
Coil 0.46 57.68 34.33 0.46 60.27 39.85
WA 0.46 64.65 48.46 0.46 67.09 52.30
BiC 0.46 62.38 43.08 0.46 65.08 50.79
FOSTER 0.46 63.38 49.42 0.46 66.49 53.21
DER 9.27 67.99 53.95 4.60 69.74 58.59
MEMO 7.14 68.10 54.23 3.62 70.20 58.49
DyTox 10.7 68.06 52.23 10.7 71.07 58.72
L2P 85.7 84.00 78.96 85.7 89.35 83.39
(a) ImageNet100 Base0 Inc5
(b) ImageNet100 Base0 Inc10
(c) ImageNet100 Base50 Inc10
(d) ImageNet100 Base50 Inc25
Figure 7: Incremental accuracy of different methods on ImageNet100. Legends are shown in (d).
Dataset Split: Following the protocol defined in [32], all classes are first shuffled by Numpy random seed $1993$. Afterward, there are two different ways to split the classes into incremental stages:
  • Train from scratch (TFS): splits all the classes equally into each incremental stage. For example, if there are $B$ stages and $C$ classes in total, each incremental task contains C/B{C}/{B} classes for training.
  • Train from half (TFH): splits half of the total classes as the first incremental task and equally assigns the rest classes into the following stages. Specifically, it assigns $C/2$ classes to the first task and C/2(B1)C/2(B-1) classes to the rest tasks.
Both of these settings are widely adopted in the current CIL community [93, 87]. Hence, to unify these settings, we use ‘Base-$m$, Inc-$n$’ to denote these settings where $m$ stands for the number of classes in the first stage, and $n$ stands for the number of classes in each incremental task. $m=0$ stands for the ‘Train from scratch’ protocol. We use the same training splits for every compared method for a fair comparison. The testing set is the same as the original one for holistic evaluation.
4.1.2 Evaluation Metrics
There are several metrics to evaluate the CIL model. We denote the Top-1 accuracy after the $b$-th task as $\mathcal{A}_{b}$, and higher $\mathcal{A}_{b}$ indicates a better prediction accuracy. Since the CIL model is continually updated, the accuracy often decays with more tasks incorporated. Hence, the accuracy after the last stage ($\mathcal{A}_{B}$) is a proper metric to measure the overall accuracy among all classes.
On the other hand, only comparing the final accuracy ignores the performance evolution along the learning trajectory. Hence, another metric denoted as ‘average accuracy’ considers the performance of all incremental stages: $\bar{\mathcal{A}}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}_{b}$. A higher average accuracy denotes a stronger performance along the incremental stages. Apart from Top-1 accuracy, some works address Top-5 accuracy for the ImageNet dataset. In the supplementary, we also report the Top-5 accuracy of these comparisons.
(a) ImageNet1000 Base0 Inc100
(b) ImageNet1000 Base500 Inc100
Figure 8: Incremental accuracy of different methods on ImageNet1000. Legends are shown in (b).
(a) Finetune
(b) EWC
(c) LwF
(d) Replay
(e) iCaRL
(f) BiC
(g) DER
(h) FOSTER
(i) MEMO
(j) Dytox
Figure 9: Confusion matrix of different methods on CIFAR100 Base0 Inc20 setting after the last incremental stage. The dotted red line splits the last 20 classes from the 80 classes learned in former tasks.
4.1.3 Implementation Details
Selected methods: In the comparison, we aim to contain all kinds of methods listed in the taxonomy of Table I. We systematically choose 16 methods, including: Replay [196], RMM [197] (data replay), GEM [59] (data regularization), EWC [77] (parameter regularization), FOSTER [21], MEMO [22], DER [20], DyTox [23], L2P [72] (dynamic networks), LwF [85], iCaRL [32], PODNET [98], Coil [92] (knowledge distillation), WA [115], BiC [87] (model rectify). We also report the baseline method ‘finetune,’ which updates the model with Eq. 4.
Among these methods, RMM is a specific technique to efficiently organize the memory budget, which can be orthogonally combined with other methods, and we combine it with FOSTER, denoted as FOSTER+RMM. L2P requires pre-trained ViT as the backbone model, while others are trained from scratch.
Training details: We implement the above methods with PyTorch [198] and PyCIL [199]. Specifically, we use the same network backbone for all CNN-based compared methods, i.e., ResNet32 [200] for CIFAR100 and ResNet18 for ImageNet. We use SGD with an initial learning rate of 0.1 and momentum of 0.9. The training epoch is set to 170 for all datasets with a batch size of 128. The learning rate suffers a decay of $0.1$ at 80 and 120 epochs. For ViT-based methods like DyTox and L2P, we follow the original implementation and use ConViT [201] for DyTox and pre-trained ViT-B/16 [24] for L2P. The optimization parameters of them are set according to the original paper since ViT has a different optimization preference to CNN. We follow the original paper to set the algorithm-specific parameters, e.g., splitting 10% exemplars from the exemplar set as validation for BiC, setting the temperature $\tau$ to 5 and using a 10 epochs warm-up for DER, using $\ell_{2}$ norm to normalize the fully-connected layers in WA.
Research has shown that a good starting point (i.e., the first stage accuracy $\mathcal{A}_{1}$) implies better transferability and will suffer less forgetting [112]. It must be noted that the performance of the first stage should be eliminated in order not to affect the forgetting evaluation. Hence, to make the methods share the same starting point, we utilize the same training strategy, data augmentation, and hyperparameters. We use basic data augmentation, e.g., random crop, horizontal flip, and color jitter for CIFAR100 and ImageNet.
It must be noted that Finetune, EWC, LwF, and L2P are exemplar-free methods, and we do not use any exemplar set for them. For other methods, we follow the benchmark setting to set the number of exemplars to 2,000 for CIFAR100 and ImageNet100 and 20,000 for ImageNet1000. These exemplars are equally sampled from each seen class via the herding [118] algorithm, as introduced in Definition 2.
4.2 Comparison on Small-Scale Datasets
In this section, we evaluate different methods on the small-scale dataset, i.e., CIFAR100. We compare these methods under both the TFS and TFH setting and design four data splits (two more splits are reported in supplementary). We report the incremental performance of different methods in Figure 6 and summarize the average and final performance, the number of parameters in Table II.
As we can infer from these figures, finetune shows the worst performance among all settings, verifying the fact that the model will suffer forgetting when sequentially learning new concepts. Regularizing the parameters, i.e., EWC, shows a negligible improvement over finetune. By contrast, LwF adds knowledge distillation loss to resist forgetting, which substantially improves the performance. When the exemplar set is available, directly replaying them during model updating can further enhance the performance by a substantial margin. Coil and iCaRL combine the exemplar repay and knowledge distillation and further obtain a performance boost than the vanilla replay. Model rectify methods, i.e., BiC and WA, rectify the bias in iCaRL and further improve the accuracy. However, recent methods based on dynamic networks (i.e., DER, FOSTER, MEMO, and DyTox) show competitive results with the help of multiple backbones. It indicates that saving more backbones can substantially help the model overcome forgetting. When it comes to pre-trained models, L2P obtains the best performance among all methods. However, other methods are trained from scratch, while L2P relies on the ViT pre-trained on ImageNet-21K, making it unfair to directly compare these two lines of methods.
On the other hand, we can infer from different settings that TFH requires more stability than TFS. Since the evaluation is based on the accuracy among all classes, remembering old classes becomes more critical when there is a large group of base classes.
(a) CIFAR100 Base0 Inc5
(b) CIFAR100 Base0 Inc10
(c) ImageNet100 Base50 Inc5
(d) ImageNet100 Base0 Inc10
Figure 10: Incremental accuracy of different methods with aligned memory cost. Legends are shown in (a) and (d).
4.3 Comparison on Large-Scale Datasets
In this section, we evaluate different methods on the large-scale dataset, i.e., ImageNet100 and ImageNet1000. We compare these methods under the TFS and TFH settings and design four data splits for ImageNet100 (two more splits are reported in supplementary) and two for ImageNet1000. We report the incremental performance (top-1 accuracy) of different methods in Figure 7 and Figure 8. We report the top-5 accuracy and summarize the average and final performance, the number of parameters in the supplementary. Note that GEM requires saving a large-scale matrix for solving the QP problem, which cannot be conducted with the ImageNet dataset. On the other hand, since the ViT in L2P is pre-trained on ImageNet-21K, incrementally training it on ImageNet is meaningless, and we do not report its results.
As we can infer from these figures, most methods share the same trend as CIFAR100. Replay acts as a strong baseline in both small-scale and large-scale inputs, verifying the effectiveness of exemplars in incremental learning. Dynamic networks consistently show the best performance among all settings, outperforming other methods by a substantial margin in the benchmark comparison.
(a) Memory size of Figure 6(b)
(b) Memory size of Figure 10(b)
Figure 11: Memory size of different comparison protocols. Dark bars denote the budget for saving exemplars, and red bars represent the budget for keeping the model in memory. Different methods should be aligned to the same budget for a fair comparison, as shown in (b).
4.4 Visualization of Confusion Matrix
Incremental models are biased toward the latest task, triggering the development of model rectify-based methods. In this section, we choose a typical phenomenon to show the model bias and how the CIL models rectify them, i.e., confusion matrix. We also visualize the weight norms of different methods in the supplementary.
Specifically, we show the confusion matrix of 10 typical methods in Figure 9 and report the full results in supplementary. The matrix is drawn under CIFAR100 Base0 Inc20 setting after the last incremental task. We can infer that finetune and EWC forget most knowledge from former classes, while LwF shows less forgetting with knowledge distillation. When equipping the model with exemplars, the diagonal elements become brighter, implying that forgetting is alleviated. Dynamic networks still show the most competitive performance in these visualizations, indicating that the former backbones help the model retrieve old features and resist catastrophic forgetting in the incremental learning process.
Table III: Average and last accuracy performance comparison on CIFAR100 with aligned memory cost. ‘#P’ represents the number of parameters (million). $\mathcal{E}$’ denotes the number of exemplars, and ‘MS’ denotes the memory size (MB). The full table is reported in the supplementary.
Method CIFAR100 Base0 Inc10
#P $\mathcal{E}$ MS $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$}
GEM 0.46 7431 23.5 27.03 10.72
Replay 0.46 7431 23.5 69.97 55.61
iCaRL 0.46 7431 23.5 70.94 58.52
PODNet 0.46 7431 23.5 60.80 45.38
Coil 0.46 7431 23.5 70.69 54.40
WA 0.46 7431 23.5 69.55 59.26
BiC 0.46 7431 23.5 70.69 59.60
FOSTER 0.46 7431 23.5 72.28 59.39
DER 4.60 2000 23.5 71.47 60.26
MEMO 3.62 3312 23.5 72.37 61.98
4.5 Memory-Aligned Comparison
As we can infer from former experimental evaluations, dynamic networks show the best performance among all methods. However, are these methods fairly compared? We argue that these dynamic networks implicitly introduce an extra memory budget, namely model buffer for keeping old models. The additional buffer results in an unfair comparison to those methods without storing models. Taking Table II for an example, the number of parameters in DER is ten times that of iCaRL in the CIFAR100 Base0 Inc10 setting. We visualize the memory cost of different methods in Figure 11(a) and find that dynamic networks obtain better performance at the expense of more memory budgets. It makes directly comparing different kinds of methods unfair.
In this section, we follow [22] and aim for a fair comparison among different methods with different memory budgets. For those methods with different memory costs, we need to align the performance measure at the same memory scale for a fair comparison. For example, a ResNet32 model costs $463,504$ parameters (float), while a CIFAR image requires $3\times 32\times 32$ integer numbers (int). Hence, the budget for saving a backbone is equal to saving $463,504$ floats $\times 4$ bytes/float $\div(3\times 32\times 32)$ bytes/image $\approx 603$ instances for CIFAR. A fair comparison between different methods can be made by equipping the other methods with more exemplars. Since DER requires saving all the backbones from history, we align the memory cost of other methods to DER with extra exemplars. We visualize the memory budget of different methods under the current protocol in Figure 11(b), where the total budgets of different methods are aligned to the same scale.
We report the results under the current protocol in Figure 10 and the detailed performance in Table III. Since Finetune, LwF, and EWC cannot be combined with the exemplar set, we do not report the results of these methods. Similarly, DyTox and L2P utilize different kinds of backbones, and we do not report their results. As we can infer from these results, the gap between dynamic networks and other methods is no longer large under fair comparison. For example, DER outperforms iCaRL by $9.07$ in terms of the final accuracy in the benchmark comparison of CIFAR100 Base0 Inc10, while the gap decreases to $1.74$ under the current protocol.
(a) CIFAR100 Base0 Inc10
(b) ImageNet100 Base50 Inc5
Figure 12: Average performance-memory curve of different methods with different datasets. Dynamic networks perform better with large budgets, while other methods dominate small ones.
4.6 Memory-Agnostic Measure
Section 4.5 enables fair comparison among different methods by aligning the memory cost. However, the comparison is made by aligning the memory cost of other methods to DER. In contrast, open-world applications require conducting class-incremental learning in various scenarios, i.e., high-performance computers and edge devices are both essential. Hence, it requires a memory-agnostic measure for CIL to measure the model’s extendability given any memory budget.
To this end, we can set several ‘comparison budgets’ and align the memory cost of different methods to them. The budget list starts from a small value and incrementally enlarges, containing the requirement of different scale models. In this setting, we set the budget of the start point to a single backbone and gradually increase it to the budget cost of DER. For those algorithms without extra model storage, we equip more exemplars to meet the selected budget. However, when the selected budget is smaller than the required size, dynamic networks (i.e., DER and MEMO) cannot be deployed with the benchmark backbone. Therefore, we choose smaller backbones with fewer parameters for alignment. Hence, we can measure the performance of different methods at different memory scales and draw the performance-memory curve, as shown in Figure 12. The X-coordinate corresponds to the memory cost, and the Y-coordinate indicates the average performance $\bar{\mathcal{A}}$ (or last performance $\mathcal{A}_{B}$}). We suggest the area under the performance-memory curve (AUC) since the curve of each method indicates the dynamic ability with the change of model size. We calculate AUC-A and AUC-L, standing for the AUC under the average performance-memory and last performance-memory curves. The results are reported in Table IV.
As we can infer from these results, there are two main conclusions. Firstly, there exists an intersection between dynamic networks and other methods, where other methods perform better given a small memory size, while dynamic networks perform better given a large memory size. In other words, there is no free lunch in CIL, and different kinds of methods have their dominant domains. Secondly, the AUC measure provides a holistic way to measure the extendability of different methods given various memory budgets. We find that FOSTER and MEMO show competitive results regarding the AUC-A/L measure, implying their stronger extendability. It would be interesting to introduce it as the performance measure and design CIL methods for real-world applications. We report implementation details and the performance of each budget point in the supplementary.
4.7 Discussions about the Comparison
With the above experimental evaluations from three aspects, we have the following conclusions:
  • Equipping the model with exemplars is a simple and effective way to resist forgetting in CIL models.
  • Knowledge distillation performs better than parameter regularization in resisting forgetting with the same cost.
  • Model rectification can further boost the performance of other CIL models in a plug-and-play manner.
  • Pre-trained models can ease the burden of incremental learning and show very strong performance. However, since the features of pre-trained models are already available and do not need to be incrementally learned, comparing pre-trained models to other methods may be unfair.
  • Dynamic networks show the best performance in the evaluation at the cost of extra memory budgets. However, when changing the extra budgets into equal size of exemplars, the performance gap becomes smaller.
  • AUC-A and AUC-L provide a way to evaluate different CIL methods in a memory-agnostic manner, which can help select the method with extendability given any budget.
Table IV: Memory-agnostic performance measures for CIL. AUC depicts the dynamic ability with the change of memory size.
Method CIFAR100 Base0 Inc10 ImageNet100 Base50 Inc5
AUC-A AUC-L AUC-A AUC-L
GEM 4.31 1.70 - -
Replay 10.49 8.02 553.6 470.1
iCaRL 10.81 8.64 607.1 527.5
PODNet 9.42 6.80 701.8 624.9
Coil 10.60 7.82 601.9 486.5
WA 10.80 8.92 666.0 581.7
BiC 10.73 8.30 592.7 474.2
FOSTER 11.12 9.03 638.7 566.3
DER 10.74 8.95 699.0 639.1
MEMO 10.85 9.03 713.0 654.6
Figure 13: Future directions of class-incremental learning. Class-incremental learning often emerges with other open-world problems, e.g., few-shot, long-tailed, multi-modal, weak supervision.
5 Future Directions
In this section, we discuss the possible future directions of the class-incremental learning field.
CIL with Complex Inputs: In the real world, data is often with complex format [202, 203], e.g., few-shot [204], imbalanced [136], with noisy label [205], weak supervised [206], multi modal [207], concept drift [13], novel classes [208, 209] etc. Incremental learning methods should be able to handle these real-world scenarios for better generalizability. Figure 13 depicts possible future applications of class-incremental learning.
Specifically, [104, 106, 210, 211] address the few-shot class-incremental learning problem, where the model is required to be adapted to incoming few-shot classes. [212, 213, 214] propose to tackle the long-tailed class-incremental learning problem, where the head classes are easy to collect with adequate instances while tail classes are scarce. Recently, with the prosperity of CLIP [147], incrementally training the language vision models to handle multi-modal data streams [188, 189] is becoming popular. Since the labeling cost is always expensive in the real world, there are some works addressing training CIL models in a semi-supervised [215] or unsupervised manner [171]. [33] proposes a unified framework to handle CIL with concept drift. If the test dataset contains instances from unknown classes, open-set recognition [208] and novel class discovery [216, 217] can equip the model with the detection ability, which refers to the open-world recognition problem [218]. Besides, test-time adaptation [219] enables the model to handle these drifts on-the-fly. Lastly, real-world data may emerge hierarchically, and the labels could evolve from coarse-grained to fine-grained [220]. Learning incremental models with refined concepts [221, 222, 223, 224, 225] is also vital to building real-world learning systems.
CIL with General Data Stream: Current CIL methods have a set of restrictions on the data stream, e.g., saving exemplars for rehearsal, undergoing multi-epoch offline training within an incremental task, etc. Future class-incremental learning algorithms should be able to conduct fully online training [226, 227, 228, 229, 230] without requiring the task boundaries [231, 232, 233]. On the other hand, saving exemplars from the history may violate user privacy in some cases. To this end, exemplar-free CIL [58, 139, 138, 57] should be conducted to enable the model to be adapted without the help of exemplars. Lastly, most CIL methods rely on the number of base classes to define hyper-parameters in model optimization, where more base classes require larger stability and fewer base classes require larger plasticity. Therefore, designing the algorithm to handle CIL problems given any base classes [234] is also important to the real world.
CIL with Any Memory Budgets: Dynamic networks have obtained impressive performance in recent years [21, 20]. However, most of them require an extra memory budget to save the external backbones. In the real world, a good CIL algorithm should handle different budget restrictions. For example, an algorithm should be able to learn with high-performance computers or with edge devices (e.g., smartphones), and both scenarios are essential. Hence, deploying and comparing CIL models in the real world should take the memory budgets into consideration. The performance measures of AUC-A/L [22] are proper solutions to holistically compare different methods given any memory budgets. On the other hand, future CIL methods are encouraged to handle specific learning scenarios. For example, [235] addresses training CIL systems under resource-limited scenarios.
CIL with Pre-trained Models: Recently, pre-trained models have shown to work competitively with their strong transferability, especially for ViT-based methods [73, 72, 74, 75]. The pre-trained models provide generalizable features for the downstream tasks, enabling the incremental model to adjust with minimal cost [145]. CIL with pre-trained models is a proper way to handle real-world incremental applications from an excellent starting point. However, since the final target for incremental learning is to build a generalizable feature continually [236], some will argue that pre-trained models weaken the difficulty of incremental learning. From this perspective, designing proper algorithms to train the CIL model from scratch is more challenging. On the other hand, pre-trained language-vision models [147, 237] have shown powerful generalizability in recent years, and exploring the ensemble of various pre-trained models [238, 239] is also an exciting topic.
CIL with Bidirectional Compatibility: Compatibility is a design characteristic in software engineering community [240, 241, 242, 243], which was introduced to the machine learning community in [244, 245]. Backward compatibility allows for interoperability with an older legacy system. In contrast, forward compatibility allows a system to accept input intended for a later version of itself. In the incremental learning field, compatibility is also a core problem, where the new model is required to understand the features produced by the old model to classify old classes. Most methods aim to enhance backward compatibility by making the new model similar to the old one. However, FACT [106] addresses the forward compatibility in CIL, where the model should be prepared for future updates, acting like the pre-assigned interface. Furthermore, it will be interesting to address bidirectional compatibility [92] in the future.
Analyzing the Reason Behind Catastrophic Forgetting: Model rectification-based methods aim to reduce the inductive bias in the CIL model. Recently, more works have tried to analyze the reason for forgetting in CIL. [172, 246] find that batch normalization layers are biased when sequentially trained, which triggers the different activation of old and new classes. [138] finds that representations with large eigenvalues transfer better and suffer less forgetting, and [226] finds that eigenvalues can be enlarged by maximizing the mutual information between old and new features. [247] theoretically decomposes the CIL problem into within-task and task-id predictions. It further proves that good within-task and task-id predictions are necessary and sufficient for good CIL performances. [248] finds that applying data replay causes the newly added classes’ representations to overlap significantly with the previous classes, leading to highly disruptive parameter updates. [249] empirically finds that the amount of forgetting correlates with the geometrical properties of the convergent points. It would be interesting to explore more reasons for catastrophic forgetting theoretically and empirically in the future.
6 Conclusion
Real-world applications are often faced with streaming data, with which the model should be incrementally updated without catastrophic forgetting. In this paper, we provide a comprehensive survey about deep class-incremental learning by categorizing them into three categories and six subcategories taxonomically and chronologically. Additionally, we provide a holistic comparison among different methods on several publicly available datasets. With these results, we discuss the insights and summarize the common rules to inspire future research. Finally, we highlight an important factor in CIL comparison, namely memory budget, and advocate evaluating different methods holistically by emphasizing the effect of memory budgets. We provide a comprehensive evaluation of different methods given specific budgets as well as some new performance measures. We expect this survey to provide an effective way to understand current state-of-the-art and speed up the development of the class-incremental learning field.
Appendix A Supplied Results in Benchmark Comparison
In this section, we report the full results of the benchmark comparison. With a bit of redundancy, we list the incremental performance of all dataset splits. Specifically, we first report the detailed results of CIFAR100 and ImageNet100 and then report the top-5 accuracy of the ImageNet comparison.
A.1 Detailed Results of CIFAR100
We first report the results of CIFAR100 where we select six dataset splits (i.e., Base0 Inc5, Base0 Inc10, Base0 Inc20, Base50 Inc50, Base50 Inc10, Base50 Inc25). The results are shown in Figure 14, and we list the number of parameters, the average accuracy, and the last accuracy in Table V.
As we can infer from these figures, L2P consistently obtains the best performance among all settings. However, it must be noted that L2P relies on the ImageNet-21K pre-trained Vision Transformer with 85M parameters. At the same time, typical methods train ResNet32 with 0.46M parameters from scratch. Hence, comparing pre-trained model based methods to typical methods is unfair. We also notice that dynamic networks, e.g., DER and MEMO obtain better performance at the cost of more parameters. Therefore, we report Table V to point out that parameter cost is sometimes ignored in the benchmark comparison, which shall lead to unfair results. These comparison details also motivate us to design fair comparison protocols for class-incremental learning.
Table V: Average and last accuracy performance comparison on CIFAR100. ‘#P’ represents the number of parameters (million).
Method Base0 Inc5 Base0 Inc10 Base0 Inc20 Base50 Inc10
#P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$} #P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$} #P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$} #P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$}
Finetune 0.46 17.59 4.83 0.46 26.25 9.09 0.46 37.90 17.07 0.46 22.79 9.09
EWC 0.46 18.42 5.58 0.46 29.73 12.44 0.46 39.19 19.87 0.46 25.77 11.47
LwF 0.46 30.93 12.60 0.46 43.56 23.25 0.46 48.96 30.00 0.46 41.12 25.06
GEM 0.46 31.73 19.48 0.46 40.18 23.03 0.46 50.33 34.09 0.46 33.28 21.33
Replay 0.46 58.20 38.69 0.46 59.31 41.01 0.46 60.03 43.08 0.46 52.37 41.26
RMM 0.46 65.72 51.10 0.46 68.54 56.64 0.46 70.64 61.81 0.46 67.53 59.75
iCaRL 0.46 63.51 45.12 0.46 64.42 49.52 0.46 67.00 54.23 0.46 61.29 52.04
PODNet 0.46 47.88 27.99 0.46 55.22 36.78 0.46 62.96 49.08 0.46 64.45 55.21
Coil 0.46 57.68 34.33 0.46 60.27 39.85 0.46 63.33 45.54 0.46 55.71 41.24
WA 0.46 64.65 48.46 0.46 67.09 52.30 0.46 68.51 57.97 0.46 64.32 55.85
BiC 0.46 62.38 43.08 0.46 65.08 50.79 0.46 67.03 56.22 0.46 61.01 49.19
FOSTER 0.46 63.38 49.42 0.46 66.49 53.21 0.46 68.60 58.67 0.46 65.73 57.82
DER 9.27 67.99 53.95 4.60 69.74 58.59 2.30 70.82 62.40 2.76 68.24 61.94
MEMO 7.14 68.10 54.23 3.62 70.20 58.49 1.87 70.43 61.39 2.22 69.39 62.83
DyTox 10.7 68.06 52.23 10.7 71.07 58.72 10.7 73.05 64.22 10.7 69.07 60.35
L2P 85.7 84.00 78.96 85.7 89.35 83.39 85.7 90.17 86.91 85.7 85.21 79.34
A.2 Detailed Results of ImageNet100
We report the results of ImageNet100 where we select six dataset splits (i.e., Base0 Inc5, Base0 Inc10, Base0 Inc20, Base50 Inc50, Base50 Inc10, Base50 Inc25). The results are shown in Figure 15, and we list the number of parameters, the average accuracy, and the last accuracy in Table VI. The conclusions are consistent with former small-scale datasets, and dynamic networks perform better at the cost of more memory budget.
It must be noted that there are two methods not included in the ImageNet comparison, i.e., GEM and L2P. GEM requires solving the QP problem and saving a large-scale matrix, which cannot be deployed with ImageNet. Besides, L2P utilizes an ImageNet-21K pre-trained backbone for class-incremental learning [72], which overlaps with the dataset to be learned. Hence, these two methods are not included in the comparison.
(a) ImageNet100 Base0 Inc5
(b) ImageNet100 Base0 Inc10
(c) ImageNet100 Base0 Inc20
(d) ImageNet100 Base50 Inc50
(e) ImageNet100 Base50 Inc10
(f) ImageNet100 Base50 Inc25
Figure 15: Incremental top-1 accuracy of different methods on ImageNet100. Legends are shown in (f).
Table VI: Average and last top-1 accuracy performance comparison on ImageNet100. ‘#P’ represents the number of parameters (million).
Method Base0 Inc5 Base0 Inc10 Base0 Inc20 Base50 Inc10
#P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$} #P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$} #P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$} #P $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$}
Finetune 11.17 17.06 4.70 11.17 26.19 9.30 11.17 40.20 17.86 11.17 24.12 9.26
EWC 11.17 18.78 6.14 11.17 27.78 11.10 11.17 41.54 18.98 11.17 26.21 11.54
LwF 11.17 41.76 17.74 11.17 55.50 33.10 11.17 68.43 53.00 11.17 46.24 31.42
Replay 11.17 56.37 37.32 11.17 59.21 41.00 11.17 64.53 48.76 11.17 55.73 43.38
RMM 11.17 69.70 57.16 11.17 74.07 65.66 11.17 80.08 73.96 11.17 73.02 66.52
iCaRL 11.17 62.36 44.10 11.17 67.11 50.98 11.17 73.57 61.50 11.17 62.56 53.68
PODNet 11.17 53.70 33.34 11.17 64.03 45.40 11.17 71.99 58.04 11.17 73.83 62.94
Coil 11.17 56.21 34.00 11.17 61.91 41.50 11.17 69.49 53.54 11.17 59.80 43.40
WA 11.17 62.96 46.06 11.17 68.60 55.04 11.17 74.44 64.84 11.17 65.81 56.64
BiC 11.17 58.03 34.56 11.17 65.13 42.40 11.17 76.29 66.92 11.17 66.36 49.90
FOSTER 11.17 64.45 53.18 11.17 69.36 60.58 11.17 75.27 68.88 11.17 69.85 63.12
DER 223.4 73.79 63.66 111.7 77.08 66.84 55.85 78.56 72.10 67.02 77.57 71.10
MEMO 170.6 68.19 56.10 86.72 71.00 60.96 44.75 76.59 68.64 53.14 76.66 70.22
DyTox 11.00 69.57 53.82 11.00 73.40 61.78 11.00 76.81 68.78 11.00 74.65 65.76
A.3 ImageNet Top-5 Results
In this section, we report the top-5 accuracy of different methods in the ImageNet100 comparison. As shown in Figure 16, the ranking of different methods remains the same as in the top-1 accuracy comparison. We also report the top-5 accuracy in the ImageNet1000 comparison in Figure 17.
(a) ImageNet100 B0 Inc5
(b) ImageNet100 B0 Inc10
(c) ImageNet100 B0 Inc20
(d) ImageNet100 B50 Inc50
(e) ImageNet100 B50 Inc10
(f) ImageNet100 B50 Inc25
Figure 16: Incremental top-5 accuracy of different methods on ImageNet100. Legends are shown in (d).
(a) ImageNet1000 B0 Inc100
(b) ImageNet1000 B500 Inc100
Figure 17: Incremental top-5 accuracy of different methods on ImageNet1000. Legends are shown in (b).
(a) Finetune
(b) EWC
(c) LwF
(d) Replay
(e) GEM
(f) iCaRL
(g) BiC
(h) WA
(i) PODNet
(j) DER
(k) RMM+FOSTER
(l) Coil
(m) FOSTER
(n) MEMO
(o) Dytox
Figure 18: Confusion matrix of different methods on CIFAR100 Base0 Inc20 setting after the last incremental stage.
(a) Finetune
(b) EWC
(c) LwF
(d) Replay
(e) GEM
(f) iCaRL
(g) BiC
(h) WA
(i) PODNet
(j) DER
(k) RMM+FOSTER
(l) Coil
(m) FOSTER
(n) MEMO
(o) Dytox
Figure 19: Weight norm of different methods on CIFAR100 Base0 Inc20 setting. We visualize the norm after the last incremental stage.
Appendix B Supplied Visualization Results
In this section, we report the full visualization results in the CIFAR100 comparison. Specifically, we first show the confusion matrix of 15 methods in Figure 18. As we can infer from these figures, finetune and EWC forget most knowledge from former classes, while LwF shows less forgetting with knowledge distillation. When comparing LwF to EWC, we can find that knowledge distillation is a more suitable solution to resist catastrophic forgetting than parameter regularization without saving exemplars. When equipping the model with exemplars, the diagonal elements become brighter, implying that forgetting is alleviated. The performance of Replay shows more steady improvement than finetune, indicating the critical role of exemplars in CIL. However, when comparing iCaRL to GEM, we can find that data regularization shows inferior performance than knowledge distillation in the efficiency of using exemplars. Moreover, BiC and WA improve the performance of iCaRL by rectifying the bias in the model, indicating that these methods can be orthogonally equipped to other CIL methods for rectification. Lastly, dynamic networks (i.e., DER, FOSTER, MEMO, and DyTox) still show the most competitive performance in these visualizations, indicating that the former backbones help the model retrieve old features and resist forgetting.
Additionally, we visualize the weight norm of the fully-connected layers in Figure 19 to show the bias introduced by incremental learning. As we can infer from these figures, when sequentially finetuning the model, the weight norm of new classes will dominate the learning process due to data imbalance. As shown in Figure 19(a), the weight norm of finetune has a stair-step shape. Since the features are passed through the ReLU layer to avoid negative ones, the larger weight norm of new classes will result in an imbalance in the prediction results. These results are consistent with the confusion matrix in Figure 18, and we can find that most CIL methods can relieve the imbalance in the fully connected layers.
Table VII: Average and last accuracy performance comparison with aligned memory cost. ‘#P’ represents the number of parameters (million). $\mathcal{E}$’ denotes the number of exemplars, and ‘MS’ denotes the memory size (MB).
Method CIFAR100 Base0 Inc5 CIFAR100 Base0 Inc10 ImageNet100 Base0 Inc10
#P $\mathcal{E}$ MS $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$} #P $\mathcal{E}$ MS $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$} #P $\mathcal{E}$ MS $\bar{\mathcal{A}}$ $\mathcal{A}_{B}$}
GEM 0.46 13466 41.22 20.30 7.99 0.46 7431 23.5 27.03 10.72 - - - - -
Replay 0.46 13466 41.22 74.25 61.09 0.46 7431 23.5 69.97 55.61 11.17 4671 713.2 67.46 53.72
iCaRL 0.46 13466 41.22 75.17 61.88 0.46 7431 23.5 70.94 58.52 11.17 4671 713.2 67.16 51.58
PODNet 0.46 13466 41.22 59.03 42.98 0.46 7431 23.5 60.80 45.38 11.17 4671 713.2 68.43 53.44
Coil 0.46 13466 41.22 74.39 60.78 0.46 7431 23.5 70.69 54.40 11.17 4671 713.2 69.97 54.24
WA 0.46 13466 41.22 74.07 63.18 0.46 7431 23.5 69.55 59.26 11.17 4671 713.2 72.62 62.12
BiC 0.46 13466 41.22 71.27 60.98 0.46 7431 23.5 70.69 59.60 11.17 4671 713.2 72.83 61.98
FOSTER 0.46 13466 41.22 75.06 62.82 0.46 7431 23.5 72.28 59.39 11.17 4671 713.2 71.55 59.86
DER 9.27 2000 41.22 70.41 57.85 4.60 2000 23.5 71.47 60.26 111.7 2000 713.2 75.70 67.14
MEMO 7.14 4771 41.22 73.57 62.97 3.62 3312 23.5 72.37 61.98 86.72 2652 713.2 72.55 64.08
(a) Average accuracy curve of CIFAR100 Base0 Inc10.
(b) Average accuracy curve of ImageNet100 Base50 Inc5.
(c) Last accuracy curve of CIFAR100 Base0 Inc10.
(d) Last accuracy curve of ImageNet100 Base50 Inc5.
Figure 20: Performance-memory curve of different methods with different datasets.
Appendix C Supplied Results in Memory-Aligned Comparison
In the main paper, we report the incremental performance curve with aligned memory to DER. In this section, we report the detailed setting of those figures, including the number of exemplars, the number of parameters, the total memory size, and the detailed performance in Table VII. Specifically, since DER and MEMO require an extra model budget than other methods, we equip more exemplars to other methods to make sure the total memory size is aligned to the same scale.
As we can infer from the table, the gap between different methods becomes smaller when the total budget is aligned to the same scale. Specifically, the typical baseline method iCaRL shows the best average accuracy in the CIFAR100 Base0 Inc5 setting. In other words, saving exemplars can be more memory-efficient than saving backbones when fairly compared.
Appendix D Details About the AUC-A/L Measure
In this section, we report the details in the memory-agnostic measure, i.e., AUC-A/L for class-incremental learning. We first highlight the importance of this measure and then give the implementation details of the curve, including the per-node performance and memory alignment details. We will start with CIFAR100 [31] and then discuss ImageNet100 [4].
D.1 Necessity of AUC in CIL
We report the full performance-memory curve in Figure 20, including the average accuracy-memory curve and last accuracy-memory curve. In each figure, the X-coordinate indicates the memory budget of the specific model, and the Y-coordinate represents the corresponding performance of the model. To highlight the importance of the AUC measure, we begin with a simple question:
Given the same memory budget, which algorithm should I choose for class-incremental learning? iCaRL or DER?
It is a simple question that aims to explore a better algorithm with the same memory budget. However, when we compare the performance of iCaRL and DER in Figure 20(a), we find an intersection among these methods. Specifically, we find iCaRL has better accuracy when the memory size is 7.4MB, while DER works better when the memory size is 23.5MB. This results in a contradictory conclusion, and we cannot find a model that suits all memory budgets.
To this end, calculating AUC does not rely on the specific budget, and we can tell from the AUC table that DER has a better AUC performance, which means it has better expandability. In other words, accuracy can only measure the performance given a specific X-coordinate. In contrast, AUC measures the area under the incremental performance curve holistically.
D.2 Implementation Details
We give the implementation details in plotting the performance-memory curve in this section. In the following discussions, we use $\mathcal{E}$ to represent the exemplar set. $\mathcal{E}$ denotes the number of exemplars, and $S(\mathcal{E})$ represents the memory size (in MB) that saving these exemplars consume. Following the benchmark implementation in the main paper, 2,000 exemplars are saved for every method for CIFAR100 and ImageNet100. Hence, the exemplar size of each method is denoted as $\mathcal{E}$=2000+E=2000+E, where $E$ corresponds to the extra exemplars exchanged from the model size, as discussed in the main paper. We use ‘#P’ to represent the number of parameters and ‘MS’ to represent the memory budget (in MB) it costs to save this model in memory. The total memory size (i.e., numbers on the X coordinate) is the summation of exemplars and the models:
$\text{Memory Size}=\text{Model Size}+\text{Exemplar Size}\,$, (18)
which should be aligned for fair comparison as we advocated.
Specifically, we can divide the selected methods into two groups. The first group contains GEM [59], iCaRL [32], Replay [196], WA [115], PODNet [98], Coil [92], BiC [87] and FOSTER [21], whose use a single backbone for incremental learning in the inference stage. Hence, they use the same network backbone with the same model size and equal memory sizes, and we denote them as SingleNet. The second group contains DER [20] and MEMO [22], which require more memory budget to save the extra model during inference. Specifically, DER sacrifices the memory size to store the backbone from history, which consumes the largest model size. On the other hand, compared to DER, MEMO does not keep the duplicated generalized blocks from history and saves much memory size to change into exemplars.
Discussion about selected methods: In this part, we choose ten methods for the comparison, i.e., Finetune, EWC, LwF, RMM, DyTox, and L2P are not included in the evaluation. Firstly, non-exemplar-based methods (Finetune, EWC, LwF) are unsuitable for the evaluation since the extendability relies on adding extra exemplars. Secondly, RMM advocates another evaluation protocol in terms of the memory, which is incompatible with ours. Lastly, DyTox and L2P rely on the vision transformer as the backbone, which have a much larger memory scale than benchmark backbones. Directly comparing ResNet to ViT may be unfair since these backbones have different characteristics in model optimization.
How to choose the budget list? As we can infer from Figure 20, there are several selected memory sizes in the figures, which formulate the X-coordinate. Firstly, the start point corresponds to the memory size of the first group, i.e., a single backbone with $2,000$ exemplars. It is a relatively small budget, which can be seen as the budget for edge devices. Correspondingly, since DER has the largest model size among all compared methods, we set the endpoint to the memory size of DER. After selecting the start and end points, we choose several intermediate budgets to formulate the budget list. The intermediate budgets are set according to the parameter size of DER and MEMO, as shown below.
How to set the exemplar and model size? For SingleNet, we can easily extend them by adding the number of exemplars. Since the model size is fixed for them, adding exemplars enables these methods to extend to a larger scale. However, we cannot use the same backbone as SingleNet for DER and MEMO when the memory size is small. Hence, we divide the model parameters into ten equal parts (since there are ten incremental tasks in the CIFAR100 Base0 Inc10 setting) and look for a backbone with similar parameter numbers. For example, we separately use ConvNet, ResNet14, ResNet20, and ResNet26 as the backbone for these methods to match different memory scales. Specifically, we use the same backbone for DER and MEMO and another same backbone for SingleNet to keep the total memory budget at the same scale. We annotate the backbone type in the figures of configurations, e.g., ‘ResNet32/ResNet14’ means we use ResNet32 for SingleNet and ResNet14 for DER and MEMO.
In the following discussions, we give the memory figures and tables to illustrate the implementation of different methods and report their incremental performance. We start with CIFAR100 and then discuss ImageNet100.
D.2.1 CIFAR100 Implementations
There are five X coordinates in the curve of CIFAR100, e.g., $7.6$, $12.4$, $16.0$, $19.8$, and $23.5$ MB. Following, we show the detailed implementation of different methods at these scales.
(a) Memory Usage
(b) Performance Curve
Figure 21: Implementation details and performance curve of CIFAR100 when memory size=7.6 MB.
Table VIII: Numerical details when memory size=7.6 MB.
7.6MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 2000 5.85MB ResNet32 0.46M 1.76MB
DER 2096 6.14MB ConvNet2 0.38M 1.48MB
MEMO 2118 6.20MB ConvNet2 0.37M 1.42MB
CIFAR100 with 7.6 MB Memory Size: The implementations are shown in Figure 21, and the memory size ($7.6$ MB) is relatively small. Since we need to align the total budget of these methods, we are only able to use small backbones for DER and MEMO. These small backbones, i.e., ConvNet with two convolutional layers, have much fewer parameters than ResNet32 and saving 10 ConvNets matches the memory size of a single ResNet32 (1.48MB versus 1.76MB). We can infer that DER and MEMO are restricted by the learning ability of these small backbones, which perform poorly in the base session. These results are consistent with the conclusions in Figure 20 that dynamic networks are inferior to SingleNet given a small memory budget.
We report the implementation details, including the number of exemplars, size of exemplars, type of backbones, number of parameters, and size of models in Table VIII. The total memory budget, i.e., 7.6MB, is the summation of the exemplar size and model size.
(a) Memory Usage
(b) Performance Curve
Figure 22: Implementation details and performance curve of CIFAR100 when memory size=12.4 MB.
Table IX: Numerical details when memory size=12.4 MB.
12.4MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 3634 10.64MB ResNet32 0.46M 1.76MB
DER 2000 5.85MB ResNet14 1.70M 6.55MB
MEMO 2495 7.32MB ResNet14 1.33M 5.10MB
CIFAR100 with 12.4 MB Memory Size: The implementations are shown in Figure 22. By raising the total memory cost to $12.4$ MB, SingleNet can utilize the extra memory size to exchange $1634$ exemplars. At the same time, dynamic networks can switch to more powerful backbones, i.e., ResNet14, to get better representation ability. We can infer that DER and MEMO show competitive results with stronger backbones and outperform other methods in this setting. These results are consistent with the conclusions in Figure 20 that the intersection between these two groups of methods exists near the start point.
(a) Memory Usage
(b) Performance Curve
Figure 23: Implementation details and performance curve of CIFAR100 when memory size=16.0 MB.
Table X: Numerical details when memory size=16.0 MB.
16.0MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 4900 14.3MB ResNet32 0.46M 1.76MB
DER 2000 5.85MB ResNet20 2.69M 10.2MB
MEMO 2768 8.10MB ResNet20 2.1M 8.01MB
CIFAR100 with 16.0 MB Memory Size: The implementations are shown in Figure 23. By raising the total memory cost to $16.0$ MB, SingleNet can utilize the extra memory size to exchange $2900$ exemplars, and dynamic networks can switch to larger backbones to get better representation ability. We use ResNet20 for DER and MEMO in this setting. The results are consistent with the former setting, where we can see that DER and MEMO show competitive results with stronger backbones and outperform other methods.
(a) Memory Usage
(b) Performance Curve
Figure 24: Implementation details and performance curve of CIFAR100 when memory size=19.8 MB.
Table XI: Numerical details when memory size=19.8 MB.
19.8MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 6165 18.06MB ResNet32 0.46M 1.76MB
DER 2000 5.85MB ResNet26 3.60M 13.9MB
MEMO 3040 8.91MB ResNet26 2.86M 10.92MB
CIFAR100 with 19.8 MB Memory Size: The implementations are shown in Figure 24. By raising the total memory cost to $19.8$ MB, SingleNet can utilize the extra memory size to exchange $4165$ exemplars, and dynamic networks can switch to larger backbones to get better representation ability. We use ResNet26 for DER and MEMO in this setting. The results are consistent with the former setting, where we can infer that dynamic networks show competitive results with stronger backbones and outperform other methods.
(a) Memory Usage
(b) Performance Curve
Figure 25: Implementation details and performance curve of CIFAR100 when memory size=23.5 MB.
Table XII: Numerical details when memory size=23.5 MB.
23.5MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 7431 21.76MB ResNet32 0.46M 1.75MB
DER 2000 5.86MB ResNet32 4.63M 17.68MB
MEMO 3312 9.7MB ResNet32 3.62M 13.83MB
CIFAR100 with 23.5 MB Memory Size: The implementations are shown in Figure 25. By raising the total memory cost to $23.5$ MB, SingleNet can utilize the extra memory size to exchange $5431$ exemplars. In addition, dynamic networks can switch to larger backbones, i.e., ResNet32, to get better representation ability.
D.2.2 ImageNet100 Implementations
Similar to CIFAR100, we can conduct an exchange between the model and exemplars with the ImageNet100 dataset. For example, saving a ResNet18 model costs $11,176,512$ parameters (float), while saving an ImageNet image costs $3\times 224\times 224$ integer numbers (int). The budget for saving a backbone is equal to saving $11,176,512$ floats $\times 4$ bytes/float $\div(3\times 224\times 224)$ bytes/image $\approx 297$ images for ImageNet. We conduct the experiment with the ImageNet100 Base50 Inc5 setting, as discussed in the main paper.
There are six X coordinates in the curve of ImageNet100, e.g., $329$, $493$, $755$, $872$, $1180$ and $1273$ MB. Following, we show the detailed implementation of different methods at these scales.
(a) Memory Usage
(b) Performance Curve
Figure 26: Implementation details and performance curve of ImageNet100 when memory size=329 MB.
Table XIII: Numerical details when memory size=329 MB.
329MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 2000 287MB ResNet18 11.17M 42.6MB
DER 2032 291MB ConvNet4 9.96M 38.0MB
MEMO 2115 303MB ConvNet4 6.81M 26.0MB
ImageNet100 with 329 MB Memory Size: The implementations are shown in Figure 26. $329$ MB is a relatively small memory size. Since we need to align the total budget of these methods, we are only able to use small backbones for DER and MEMO. These small backbones, i.e., ConvNet with four convolutional layers, have much fewer parameters than ResNet18 and saving 10 ConvNets matches the memory size of a single ResNet18. We can infer from the figure that DER and MEMO are restricted by the learning ability of the inferior backbones, which perform poorly in the base task.
(a) Memory Usage
(b) Performance Curve
Figure 27: Implementation details and performance curve ImageNet100 when memory size=493 MB.
Table XIV: Numerical details when memory size=493 MB.
493MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 3136 450MB ResNet18 11.17M 42.6MB
DER 2000 287MB ResNet10 53.96M 205MB
MEMO 2327 334MB ResNet10 41.63M 158MB
ImageNet100 with 493 MB Memory Size: The implementations are shown in Figure 27. By raising the total memory cost to $493$ MB, SingleNet can utilize the extra memory size to exchange $1136$ exemplars. In addition, dynamic networks can switch to larger backbones, i.e., ResNet10, to get better representation ability. We can infer from the figure that dynamic networks show competitive results with stronger backbones and outperform other methods in this setting.
(a) Memory Usage
(b) Performance Curve
Figure 28: Implementation details and performance curve ImageNet100 when memory size=755 MB.
Table XV: Numerical details when memory size=755 MB.
755MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 4970 713MB ResNet18 11.17M 42.6MB
DER 2000 287MB ResNet18 122.9M 468MB
MEMO 2739 393MB ResNet18 95.11M 362MB
ImageNet100 with 755 MB Memory Size: The implementations are shown in Figure 28. By raising the total memory cost to $755$ MB, SingleNet can utilize the extra memory size to exchange $2970$ exemplars. At the same time, dynamic networks can switch to larger backbones, i.e., ResNet18, to get better representation ability.
(a) Memory Usage
(b) Performance Curve
Figure 29: Implementation details and performance curve ImageNet100 when memory size=872 MB.
Table XVI: Numerical details when memory size=872 MB.
872MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 5779 829MB ResNet18 11.17M 42.6MB
DER 2000 287MB ResNet26 153.4M 585.2MB
MEMO 2915 417MB ResNet26 119.0M 453MB
ImageNet100 with 872 MB Memory Size: The implementations are shown in Figure 29. By raising the total memory cost to $872$ MB, SingleNet can utilize the extra memory size to exchange $3779$ exemplars. In addition, dynamic networks can switch to larger backbones, i.e., ResNet26 for better representation ability.
(a) Memory Usage
(b) Performance Curve
Figure 30: Implementation details and performance curve ImageNet100 when memory size=1180 MB.
Table XVII: Numerical details when memory size=1180 MB.
1180MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 7924 1137MB ResNet18 11.17M 42.6MB
DER 2000 287MB ResNet34 234.1M 893MB
MEMO 4170 598MB ResNet34 152.4M 581MB
ImageNet100 with 1180 MB Memory Size: The implementations are shown in Figure 30. By raising the total memory cost to $1180$ MB, non-dynamic networks can utilize the extra memory size to exchange $5924$ exemplars. At the same time, dynamic networks can switch to larger backbones, i.e., ResNet34, to get better representation ability.
(a) Memory Usage
(b) Performance Curve
Figure 31: Implementation details and performance curve ImageNet100 when memory size = 1273 MB.
Table XVIII: Numerical details when memory size=1273 MB.
1273MB $\mathcal{E}$ $S(\mathcal{E})$ Model Type #P Model Size
SingleNet 8574 1230MB ResNet18 11.17M 42.6MB
DER 2000 287MB ResNet50 258.6M 986MB
MEMO 4270 612MB ResNet50 173.2M 660MB
ImageNet100 with 1273 MB Memory Size: The implementations are shown in Figure 31. By raising the total memory cost to $1273$ MB, SingleNet can utilize the extra memory size to exchange $6574$ exemplars, and dynamic networks can switch to larger backbones to get better representation ability. We use ResNet50 for DER and MEMO in this setting.
Discussion about Backbones: It should be noted that ResNet18 is the benchmark backbone for ImageNet, and DER consumes about 755 MB memory budget under the benchmark setting. Hence, the last three points in the X-coordinate, i.e., 872, 1180, and 1273 MB, are larger than the benchmark setting. There are two main reasons for the budget list design. First, handling large-scale image inputs requires more convolutional layers, and it is hard to find typical models with small memory budgets. Second, we would like to investigate the performance when the model is large enough to see whether the improvement of stronger backbones will converge. The empirical results successfully verify our assumptions.